### v1

# Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution

Abstract
The success of DNNs often depends on training with large-scale datasets, but building such datasets is both expensive and challenging. Consequently, public datasets from open-source platforms like HuggingFace have become popular, posing significant risks of data poisoning attacks. Existing backdoor defenses in NLP primarily focus on identifying and removing poisoned samples; however, purifying a backdoored model with these sample-cleaning approaches typically requires expensive retraining. Therefore, we propose Greedy Module-Substitution (GMS), which identifies and substitutes “deadwood” modules (i.e., components critical to backdoor pathways) in a backdoored model to purify it. Our method relaxes the common dependency of prior model purification methods on clean datasets or clean auxiliary models. When applied to RoBERTa-large under backdoor attacks, GMS demonstrates strong effectiveness across various settings, particularly against widely recognized challenging attacks like LWS, achieving a post-purification attack success rate (ASR) of 9.7% on SST-2 compared to 58.8% for the best baseline approach.
Cut the Deadwood Out: Post-Training Model Purification
with Selective Module Substitution
Yao Tong1thanks: Equal contribution., Weijun Li2∗, Xuanli He3, Haolan Zhan4, Qiongkai Xu2thanks: The corresponding author. 1 National University of Singapore, 2 Macquarie University 3 University College London, 4 Monash University tongyao@u.nus.edu, weijun.li1@hdr.mq.edu.au xuanli.he@ucl.ac.uk, haolan.zhan@monash.edu, qiongkai.xu@mq.edu.au
1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various domains Krizhevsky et al. (2012); Bowman et al. (2016), largely driven by training on large-scale datasets. Given the high cost of collecting and labeling data from scratch, incorporating public datasets from open-source platforms like HuggingFace in training has become a common practice, particularly in NLP Howard and Ruder (2018); Ding et al. (2023). However, this growing reliance on open-source datasets from unaudited platforms raises significant security concerns Yang et al. (2021a); Zhang et al. (2021), especially for commercial use. Malicious actors can exploit this ecosystem by publishing poisoned datasets Xu et al. (2021) to embed backdoors into models for potential benefit. These backdoors allow attackers to manipulate models, causing targeted misbehavior when triggered by specific inputs while operating normally on clean inputs. As a result, they often remain stealthy until their triggers are activated.
Figure 1: Parameter-level editing (left) prunes parameters "most relevant" to backdoor features but leaves many backdoor-related parameters intact within each module, allowing backdoor pathways to potentially persist. In contrast, our module-level substitution (right) effectively disrupts these pathways by completely replacing backdoor-connected modules with counterparts from a proxy model (not required to be clean).
Existing defense strategies against backdoor attacks in NLP primarily focus on backdoor samples detection, such as detecting and removing poisoned samples leveraging their distinct characteristics compared to clean ones Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2022, 2023); Xu et al. (2024). However, their effectiveness is often suboptimal Min et al. (2024), especially as backdoor triggers evolve to be stealthier and harder to detect or reverse Qi et al. (2021b). Moreover, while these methods follow a fine-tuning or adversarial training for unlearning Wei et al. (2023) to sanitize poisoned models, they cannot effectively purify a backdoored model without incurring a significant cost of retraining Wu et al. (2022). Then a practical question arises: How can we cost-efficiently eliminate the backdoor features from an existing model after users reported some compromise?
While only a few model purification defenses in NLP exist, they predominantly focus on parameter-level editing. For instances, Fine-mixing Zhang et al. (2022) and Fine-purification Zhang et al. (2023) mix a clean pre-trained language model (PLM) with a backdoored model and rely on fine-tuning with a clean dataset for purification. Similarly, Zhao et al. (2024) prunes attention heads potentially affected by poisoned data and fine-tunes the pruned model using a clean dataset. However, recent studies have observed that backdoor features are encoded as complex pathways in victim models, with significant overlap between backdoor and clean neurons at the module level Jebreel et al. (2023); Li et al. (2024). As illustrated in Figure 1 (Left), parameter-level editing will leave hidden vulnerabilities in the non-pruned connections, thus making the defense effectiveness heavily relies on fine-tuning with a clean dataset to restore clean features Liu et al. (2018), an assumption that is both strong and rarely feasible in practice.
Here, we develop a method that effectively disrupts the model’s backdoor pathways without relying on the assumption of access to a clean dataset. Note that the abundance of public platforms for open-source models (e.g., HuggingFace) offers semantically similar yet independent models for any task domain Semwal et al. (2018). Therefore, we propose a simple yet effective module-level purification defense, Greedy Module-Substitution (GMS), which identifies critical modules for backdoor propagation and purifies them by replacing the parameters with those of a proxy model. While this approach eliminates the need for a clean dataset to restore semantic knowledge, the proxy model is not required to be clean either, since the pathways of two different backdoored models are less likely to be connected (empirically verified in Table 5). The primary contributions of this work are as follows.
  1. 1.
    We propose the first efficient (quadratic time complexity) module-level backdoor purification defense (GMS), which relaxes the strong reliance on clean auxiliary models and clean datasets required by prior defenses, as shown in Table 5 and Table 6.
  2. 2.
    Our method significantly outperforms baseline approaches against challenging backdoor attacks like LWS and HiddenKiller, e.g., on SST-2 victim model, GMS reduces the ASR on LWS to 9.7%, compared to 12.8% for the benign model and 58.8% for the best baseline.
  3. 3.
    Our GMS purification strategy exhibits strong transferability across different attacks. Even the strategy with few module substitutions can effectively defend against other backdoor attacks, making it suitable when the victim dataset is inaccessible (Section 5.3 and Table 7).
2 Related works
Backdoor attack
Backdoor attacks aim to manipulate models to behave normally on benign inputs while exhibiting attacker-controlled behavior when specific triggers are present. They can be categorized into two series Wu et al. (2022): (1) Data-poisoning attacks, exemplified by BadNets Gu et al. (2017), involve tampering with a subset of training data by adding triggers and altering labels, making them practical in real-world scenarios with minimal attacker assumptions Turner et al. (2019); Carlini and Terzis (2022); Goldblum et al. (2023); Li et al. (2021c). Initially explored in image classification, backdoor attacks have since raised significant security concerns in NLP Dai et al. (2019); Qi et al. (2021d, c) with triggers range from misspelled words Chen et al. (2021), rare words Yang et al. (2021a); Kurita et al. (2020) and syntactic structures Qi et al. (2021c) to text styles Pan et al. (2022), posing serious challenges for detection. (2) Training-control attacks, on the other hand, assume complete control over the training process and data Nguyen and Tran (2021); Kurita et al. (2020), with advances such as layer-wise weight poisoning ensuring backdoor persistence even after fine-tuning Li et al. (2021a). In this work, we focus on data-poisoning backdoor attacks in NLP due to their widespread applicability and practical implications.
Backdoor defense
Backdoor defense approaches can be categorized according to their strategies: (1) Backdoor sample detection focuses on identifying poisoned samples in the dataset. The defense approaches usually include detecting and removing backdoor samples, leveraging the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2019, 2022, 2023); Xu et al. (2024). These strategies are typically applied during the training phase or involve fine-tuning after removing backdoor samples. (2) Backdoor model purification: These methods aim to eliminate backdoor features from a well-trained model. Common approaches involve merging parameters with other proxy models Arora et al. (2024); Chen et al. (2024a), or fine-tuning or pruning using a limited amount of clean data Wu and Wang (2021); Min et al. (2023); Liu et al. (2018); Zhao et al. (2024); Zhu et al. (2023). These techniques not only reduce training costs but also achieve state-of-the-art defense performance Zhu et al. (2023); Zhao et al. (2024); Arora et al. (2024). Our work advances the second line of research by minimizing the number of auxiliary models to one and reducing dependence on the clean dataset.
3 Preliminary
Modern neural network (NN) architectures can be viewed as layer-wise compositions of functional blocks, where each block serves as a reusable unit performing a well-defined computational operation. Hereby, we define two aspects to systematically describe generic NN architectures:
  1. 1.
    Layer Set ($\mathcal{L}$) is the set of block indices, capturing the architecture’s depth and structure.
  2. 2.
    Module Set ($\mathcal{M}$) is the set of representative functional modules within each block.
Transformer blocks
In this paper, we use Transformer-based architecture Vaswani et al. (2017) as the testbed, given its widespread attention and implementation in NLP.
For each Transformer block, the input will be processed through Attention and Feed-Forward Network (FFN) modules. Attention modules contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively. The following FFN contains two components: $W_{F}$ for the forward layer and $W_{P}$ for the projection layer, connected by a non-linear activation function.
In total, each Transformer block contains six key functional modules with parameters $W_{Q},W_{K},W_{V},W_{O},W_{F},$ and $W_{P}$. For simplicity and consistency, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper. Accordingly, we define the module set $\mathcal{M}$ for Transformer models as $\{Q,K,V,O,F,P\}$ and layer set =$\mathcal{L}=\{1,2,\ldots,L\}$.
Evaluation metrics
Following previous literature (Qi et al., 2020), we adopt two metrics: (1) Clean Accuracy (CACC) measures the prediction accuracy on clean samples, and (2) Attack Success Rate (ASR) measures the attack accuracy on poisoned test set. A higher CACC reflects better utility, while lower ASR indicates stronger defense.
Figure 2: The pipeline of our method. Step 1: Extract two small proxy datasets from the suspected dataset $\mathcal{D}$ used for computing score in Equation 1. Step 2: Iteratively update a substitution matrix $S$ to greedily maximize the trade-off score between backdoor removal and utility preservation. Step 3: Return purified model $M_{pure}$ corresponding to the best substitution matrix $S_{best}$. For more details, refer to Algorithm 1.
4 Methodology
In Section 4.1, we introduce the problem setting, and in Section 4.2 we present our methods.
4.1 Problem setting
Threat Model
We consider the scenario of data poisoning-based backdoor attacks, where an attacker releases a poisoned dataset $\mathcal{D}$ to embed backdoors into any model trained on this dataset. Starting with a clean dataset $\mathcal{D}_{\text{cl}}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ denote the input and label spaces, respectively, the attacker selects a subset $\mathcal{D}_{\text{sub}}\subset\mathcal{D}_{\text{cl}}$, and then modified each sample $(x,y)\in\mathcal{D}_{\text{sub}}$ into a poisoned sample $(\tilde{x},y_{t})$ where $\tilde{x}=g(x,\Delta)$, the pre-defined trigger is $\Delta$, $g$ is a generation function, and $y_{t}$ is the target label. The poisoned subset is $\mathcal{D}_{\text{poi}}=\{(\tilde{x},y_{t})\mid(x,y)\in\mathcal{D}_{\text{sub}}\}$. Finally, the attackers publish the poisoned dataset constructed as $\mathcal{D}=(\mathcal{D}_{\text{cl}}\setminus\mathcal{D}_{\text{sub}})\cup\mathcal{D}_{\text{poi}}$.
Defense Scenario
The defender (e.g., a model vendor) has already trained a (potentially backdoored) victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$. While the defender has access to $\mathcal{D}$, they are unaware of all possible backdoor patterns. [^1] As shown in Section 4.2.2, the full training dataset is not strictly required; a small subset of poisoned data reported by users or automatically detected by algorithms is sufficient for applying our defense. The defender has white-box access to a trained victim model $M_{v}$ and a proxy model $M_{p}$ trained on similar downstream tasks.
4.2 Our method
We aim to disrupt the backdoor connection in victim model $M_{v}$ while preserving the knowledge required for the original task. Models trained on similar tasks tend to capture similar domain-specific knowledge Semwal et al. (2018). Hence, we replace backdoor-connected modules in the $M_{v}$ with counterparts from a proxy model $M_{p}$ that shares similar domain knowledge. Notably, $M_{p}$ does NOT need to be benign, as even a malicious proxy model lacking identical backdoor pathways will fail to restore the disrupted backdoor connection.
Our main algorithm, Greedy Module-Substitution (GMS), is detailed in Section 4.2.3. Section 4.2.1 defines the objective optimized by GMS, and Section 4.2.2 describes the method to construct the proxy datasets used in evaluation.
4.2.1 Objective
The defender’s objective is to identify a purified model $M_{pure}$ from all possible substituted models $M$ that effectively eliminates the backdoor while preserving task utility. This can be formulated as
$\operatorname*{argmax}_{M}\,(1-\alpha)\cdot\delta_{\text{asr}}(M)+\alpha\cdot(1-\delta_{\text{acc}}(M))$ (1)
where $\delta_{\text{asr}}(M)$ measures the extent of backdoor removal by quantifying the reduction in ASR and $\delta_{\text{acc}}(M)$ assesses the impact on task utility by measuring the decrease in CACC. However, since the defender lacks access to the clean dataset and poisoned test set, we adopt the method described in Section 4.2.2 to extract two proxy datasets, $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, from the suspected dataset $\mathcal{D}$ for ASR and CACC estimation.
The goal is to maximize $\delta_{\text{asr}}(M)$ while minimizing $\delta_{\text{acc}}(M)$. The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor defense and task utility. A larger $\alpha$ favors utility, while a smaller $\alpha$ prioritizes backdoor defense.
4.2.2 Proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$
In this section, we introduce the process of extracting two proxy datasets from the suspected dataset $\mathcal{D}$: a proxy clean set $\mathcal{D}_{clean}$, which is identified with high probability as clean (i.e., they are likely to be clean though not guaranteed to be free from poisoning), and a proxy poison set $\mathcal{D}_{poison}$, which is identified with high confidence as being poisoned. Both proxy datasets are not required to be purely clean or poisoned. When the defender has no prior knowledge of the poisoned samples, various methods can be used to construct proxy datasets, such as sample diagnosis Gao et al. (2019); Huang et al. (2022); Guo et al. (2023) and trigger inversion Wang et al. (2019, 2022, 2023); Xu et al. (2024). If additional prior knowledge, such as a user-reported clean dataset, is available (as tested in Table 6), the performance of our method is expected to be improved. For simplicity, in this paper, we use random sampling to construct the proxy clean set $\mathcal{D}_{clean}$, based on the practical assumption that the poison samples will not dominate a training set. To construct the proxy poison set $\mathcal{D}_{poison}$, we adopt an existing backdoor sample detection method from SEEP He et al. (2024), which leverages an observation that backdoored models tend to exhibit over-confidence on poisoned samples Li et al. (2021b). SEEP utilizes training dynamics to identify anomalous samples in the training corpus, which are proven to have a high probability of being backdoor samples. In our example, we only focus on the most confident samples extracted (i.e., seed set). In Section 5.2, we show that extracting only 1% of the original poisoned data in $\mathcal{D}_{poison}$ and allowing even 20% poisoned data in $\mathcal{D}_{clean}$ are sufficient to launch an effective defense using our Greedy-Module-Substitution (GMS).
4.2.3 Greedy search for purified model $M_{pure}$
Using proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$, we can estimate ASR and ACC to quantify the level of backdoor defense and task utility for a model $M$, we define the following two metrics:
$\delta_{\text{asr}}(M)$ $=s_{asr}(M_{v})-s_{asr}(M),$ (2)
$\delta_{\text{acc}}(M)$ $=s_{acc}(M_{v})-s_{acc}(M).$ (3)
Here, $s_{acc}(M)$ represents the CACC of $M$ on $\mathcal{D}_{clean}$, while $s_{asr}(M)$ denotes its ASR on $\mathcal{D}_{poison}$. Similarly, $s_{acc}(M_{v})$ and $s_{asr}(M_{v})$ are the ASR and CACC for the original victim model $M_{v}$. In this context, $\delta_{\text{asr}}(M)$ quantifies the reduction in ASR achieved by a purified model $M$, while $\delta_{\text{acc}}(M)$ reflects the drop in task utility. Plugging them back into the objective (Equation 1) gives us an score. [^2] An inaccurate $\mathcal{D}_{poison}$ may lead to a smaller $\delta_{\text{asr}}$, while an impure $\mathcal{D}_{clean}$ might result in a larger $\delta_{\text{acc}}(M)$. However, since the metrics are always calculated on fixed proxy datasets, these effects can be controlled by adjusting $\alpha$.
While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\cdot 12}$ for a BERT-base (Devlin et al., 2019)). Thus, we propose a greedy algorithm that transforms the problem into a feasible search by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
Algorithm 1 Greedy-Module-Substitution (GMS)
1:  Input: Victim model $M_{v}$, proxy model $M_{p}$, module set $\mathcal{M}$, layer set $\mathcal{L}$, proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
2:  Output: Purified model $M_{pure}$
3:  Initialize $S\leftarrow\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$, $c_{best}\leftarrow-\infty$, $S_{best}\leftarrow\varnothing$
4:  $(s_{acc},s_{asr})\leftarrow$ Evaluate($M_{v}$)
5:  # Iteratively Update Selected Modules and Layers:
6:  while $|\mathcal{M}|>1$ or $|\mathcal{L}|>1$ do
7:     $c^{*}\leftarrow-\infty$, $S^{*}\leftarrow\varnothing$
8:     for each $m\in\mathcal{M}$ or $l\in\mathcal{L}$ do
9:        $S^{\prime}\leftarrow S$ with $S^{\prime}_{:,m}=\mathbf{0}_{L\times 1}$ or $S^{\prime}_{l,:}=\mathbf{0}_{1\times|\mathcal{M}|}$
10:        $M_{pure}^{\prime}\leftarrow$ Replace($M_{v}$, $M_{p}$, $S^{\prime}$)
11:        $c\leftarrow$ ComputeScore($M_{pure},s_{acc},s_{asr}$)
12:        if $c>c^{*}$  then
13:           Update $c^{*}\leftarrow c,S^{*}\leftarrow S^{\prime}$
14:        end if
15:     end for
16:     $S\leftarrow S^{*}$ and update $\mathcal{M},\mathcal{L}$ accordingly
17:     if $c^{*}>c_{best}$  then
18:        Update $c_{best}\leftarrow c^{*}$, $S_{best}\leftarrow S^{*}$
19:     end if
20:     if $c_{best}$ not been updated for $T$ iterations then
21:        break
22:     end if
23:  end while
24:  return  $M_{pure}=\text{Replace}(M_{v},M_{p},S_{best})$
Purifying $M_{v}$ by module substitution
Given a backdoored model $M_{v}$ (characterized by both high CACC on a clean dataset and high ASR on a poisoned dataset), we aim to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\delta_{\text{asr}}$).
Our algorithm, as demonstrated in Figure 2 and Algorithm 1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if =$\mathcal{M}=\{O,F,P\}$ and =$\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$. For simplicity, we introduce a substitution matrix $S\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{L}|}$ to specify which modules in which layer in the victim model $M_{v}$ are selected for substitution. Note that while the shape of $S$ is determined by the initial sizes of sets $\mathcal{M}$ and $\mathcal{L}$, these two sets are iteratively updated throughout the algorithm. Each row of $S$ corresponds to a layer in the layer set $\mathcal{L}$, and each column corresponds to a module in the module set $\mathcal{M}$. The value of $S[l,m]=1$ indicates that the module $m\in\mathcal{M}$ in layer $l\in\mathcal{L}$ in the victim model $M_{v}$ is selected for substitution, and $S[l,m]=0$ otherwise. Given $S$, we can succinctly represent the parameter substitution as $S\odot M_{p}+(1-S)\odot M_{v}$, where $\odot$ denotes the element-wise product.
We start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., =$\mathcal{M}=\{K,Q,V,O,F,P\}$, =$\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$. At $t$-th iteration, the algorithm considers two types of updates to $S^{t}$: deselecting a module from the current module set $\mathcal{M}$ (affecting all selected layers) or deselecting a layer from the layer set $\mathcal{L}$ (affecting all selected modules). The update that maximizes the score in Equation 1 is applied to $S^{t+1}$. The process continues until one of two conditions is met: (1) further iterations will not yield a better strategy (controlled by stopping patience $T$); or (2) no additional layers or modules can be removed, i.e., $|\mathcal{M}|=|\mathcal{L}|=1$. Finally,the strategy $S$ that yields a purified model $M_{pure}$ with the highest score is returned. The complexity of our algorithm is a practical quadratic $O(|\mathcal{M}|^{2}+|\mathcal{L}|^{2})$.
5 Experiments
5.1 Experimental settings
Datasets.
We evaluate our method on four datasets: SST-2 (Socher et al., 2013), OLID (Zampieri et al., 2019), MNLI (Williams et al., 2018), and AGNEWS (Zhang et al., 2015). These datasets cover text classification and natural language inference (NLI) tasks, including binary and multi-class classification scenarios, and are widely used for evaluating text backdoor attacks and defenses (Qi et al., 2021a; He et al., 2023; Gupta and Krishna, 2023; Arora et al., 2024).
We download the open-source versions of these datasets provided by HuggingFace (Lhoest et al., 2021). Table 1 presents the statistics of the evaluated datasets. To train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively. For MNLI, we use a random subset of 100,000 samples to reduce overhead.
Backdoor Attacks.
We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021d), and (4) HiddenKiller (Qi et al., 2021c). The first two correspond to insertion-based methods, while the latter two involve synonym substitution and syntactic paraphrasing approaches, respectively.
Follow Arora et al. (2024), we use rare words {"cf", "mn", "bb", "tq", "mb" } as triggers for BadNets and phrases {"I watched this movie", "no cross, no crown" } for InsertSent. The poison target labels for each dataset are listed in Table 1. While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section B.2.
Dataset Classes Train Test Target Class
Clean Poison
SST-2 2 67,349 872 444 Negative (0)
OLID 2 13,240 860 240 Not offensive (1)
MNLI 3 100,000 400 285 Neutral (1)
AGNews 4 120,000 7,600 5,700 Sports (1)
Table 1: The statistics of our evaluated datasets.
Defense Baselines.
We choose six renowned defenses as our baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def. (He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024), (4) TIES (Yadav et al., 2023), (5) DARE (Yu et al., 2024), and (6) WAG (Arora et al., 2024).
Among these, ONION and Z-Def. detect and remove triggers from datasets. ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def. detects spurious correlations between tokens and labels. In contrast, (2)-(6) are model purification methods. PURE purifies the victim model using attention head pruning and normalization techniques. TIES, DARE and WAG merge the victim model with other models to sanitize it.
Evaluation Metrics.
As introduced in  Section 3, we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure the model performance. CACC is evaluated on a clean test set, where both the samples and labels are ground truth. In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
ONION 56.8 92.9 99.9 93.3 85.7 91.9 92.9 92.9
Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
PURE 0 50.9 0 50.9 0 50.9 0 50.9
TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
DARE w/ TIES 99.3 96.0 100.0 96.2 96.4 95.7 92.7 95.8
WAG 84.4 94.8 60.1 95.2 58.8 94.8 56.2 92.5
SST-2 Ours (GMS) 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
ONION 59.4 94.8 97.8 95.1 84.8 94.5 99.6 94.7
Z-Def. 1.6 95.3 0.4 95.3 97.9 96.1 100.0 95.0
PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
DARE w/ TIES 99.9 95.2 99.6 95.4 97.8 96.5 100.0 95.2
WAG 92.7 94.1 97.8 94.4 78.0 93.9 90.9 94.3
AGNews Ours (GMS) 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
Table 2: (Partial) Performance of our method on two datasets compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds. We highlight the top-2 lowest ASR results in blue cells, and the highest CACC results in bold. Results for other datasets are in Table 8.
Implementation details
Following Arora et al. (2024), we compare all methods on RoBERTa-large (Liu et al., 2019) and BERT-base-uncased (Devlin et al., 2019). All victim models are fine-tuned on poisoned datasets for three epochs using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. All defense baselines are implemented based on their respective open-source repositories (see Section A.3). The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section 4.2.2. To train proxy models for selection, we use IMDB (Zhang et al., 2015) for the SST-2 victim model, Twitter Abusive (Founta et al., 2018) for the OLID, SNLI (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
Our method uses hyperparameter $\alpha$ in Equation 1 to control the trade-off between backdoor defense and task utility. By default, $\alpha$ is set to 0.4 but can be adjusted based on priorities: smaller values (e.g., 0.1) for lower ASR and larger values (e.g., 1.0) for higher utility. The stopping patience $T=5$ is default in Algorithm 1. In Section B.3, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
5.2 Main results
Table 2 presents the performance of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds. Results on other datasets and architectures are provided in Section B.1. Across all datasets and backdoor attacks, our method consistently ranks among the top 2, regarding backdoor removal performance, with minor harm to clean accuracy. In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines—for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def. While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large. We discuss the potential reasons for this instability in Section B.1.
Dataset Model CACC $\uparrow$
SST-2 Benign Victim Model 95.9
Proxy Model (IMDB) 88.7
Purified Model (GMS) 95.8
AGNews Benign Victim Model 95.2
Proxy Model (BBCNews) 78.1
Purified Model (GMS) 95.3
Table 3: CACC of benign victim models preserved.
BadNet InsertSent LWS HiddenKiller
Dataset Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
IMDB 5.4 92.8 1.6 94.0 12.6 93.8 11.0 93.0
YELP 4.3 96.8 8.1 94.8 23.2 92.3 19.1 93.8
SST-2 AMAZON 3.8 92.9 1.1 95.8 20.1 92.6 13.5 93.8
Table 4: Proxy models trained on different datasets can all effectively purify victim models (trained on the SST-2).
Model Dataset Backdoored Method Before Substitution After Substitution
[…]
Victim SST-2 BadNet 95.6 100.0 95.3 4.5
Proxy SST-2 HiddenKiller 95.4 7.9
Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
Proxy SST-2 BadNet 95.6 17.1
Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
Proxy IMDB 87.7 8.6
Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
Proxy IMDB 90.1 100.0
Table 5: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor does not exactly match in both attack strategy and trigger.
Additionally, although merging multiple models has been shown to be an effective backdoor defense, we observed that applying all merging baselines using a single proxy model is sub-optimal, aligning with the results in Arora et al. (2024). We found these methods prioritize preserving accuracy over removing backdoors. This behavior likely arises from the design of merging mechanisms, which are intended to preserve the performance of each merged model on downstream tasks. Consequently, backdoor tasks are treated equivalently to downstream tasks, exposing a persistent vulnerability.
5.3 Ablation studies
Utility Preservation for Benign Victim Models
While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign. Does our method compromise the utility of a clean victim model? As shown in Table 3, our results effectively dispel this concern. When applying GMS to a benign model, the CACC remains unaffected.
The effect of different proxy models
Our method relies on a proxy model for substitution, raising the natural question of how sensitive it is to the selection of proxy models. We tested two scenarios: Scenario 1: We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2. As shown in Table 4, our method is largely insensitive to the choice of proxy model across all backdoor attacks. Using any proxy model, GMS effectively defends against all tested attacks. Scenario 2: What if the proxy model itself is backdoored? We tested cases where the victim model and the proxy model were trained on the same dataset with different backdoor attacks or on different datasets with the same attacks. From Table 5, we observed that as long as the backdoor in the proxy model is not identical to that in the victim model (i.e., the same backdoor strategy with the same backdoor trigger), our GMS method consistently removes the backdoor.
Performance under different knowledge priors
While our method uses random sampling and outlier detection to extract the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, we also evaluate the potential performance gain if more accurate prior knowledge were available. For example, we may receive an accurate clean dataset or an accurate label-flipped poison dataset (i.e., dirty label) from reliable users or annotators. As shown in Table 6, with or without prior knowledge of accurate clean or poisoned data, our method consistently purifies the backdoor, achieving an ASR same as that of a benign model (4.1%). With optimal priors, GMS maintains the CACC of the benign model (95.6%), compared to 93.6% in the worst case without such prior knowledge.
Knowledge Prior Merged SST-2 (BadNet) & IMDB
[…]
Random Sampling Pure Clean Outlier Detection Dirty Label
3.2 93.6
4.1 94.9
4.3 92.7
4.3 95.6
Benign SST-2 4.1 95.6
Table 6: Performance of purifying a victim model trained on SST-2 with a proxy model trained on IMDB, using proxy datasets with varying knowledge priors. The backdoor attack strategy is BadNet.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
Table 7: (Partial) The performance of transferring strategy searched based on BadNet to other attacks. Results for other datasets are in Table 12.
Transferability of substitution strategies
Interestingly, upon examining the module substitution strategies ($S_{best}$) searched by GMS for addressing various backdoor attacks (illustrated with examples in Figure 6 in Section B.5), we observe that, despite differences among the strategies, they exhibit transferability across different backdoor attacks.
In Table 7, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure 6) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm 1 using $S_{badnet}$. Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably. This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.
6 Conclusion
We propose Greedy-Module-Substitution (GMS), an efficient and effective model purification defense against data-poisoning attacks in NLP, eliminating the need for strictly clean datasets or clean auxiliary models. GMS recursively identifies modules critical to backdoor pathways and substitutes them for those from a proxy model, which is not necessarily to be clean. Extensive experiments demonstrate the effectiveness of GMS and, more excitingly, the transferability of the optimized substitution strategies.
Limitations
Our method goes one step beyond previous auxiliary-model-based defenses by allowing the proxy model to be malicious, as long as its backdoor is not identical to the victim’s backdoor (i.e., the same attack method with exactly the same trigger, which is rare in practice) as shown in Table 5. However, further relaxation of the requirements to handle cases where the proxy model is backdoored in the same way could enable even more robust and powerful defense strategies. Exploring this direction presents an important and meaningful avenue for future research.
Appendix A Experiment details
A.1 Experiment setup
We adopt the same training configuration as WAG (Arora et al., 2024) to ensure a fair comparison. This includes fine-tuning roberta-large and bert-base-uncased on poisoned datasets for three epochs, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. The batch size and maximum sequence length are set to 32 and 128, respectively.
To validate our defense, we first adapt the aforementioned backdoor methods and datasets to obtain victim models. Next, we train proxy models on homologous datasets to prepare for future merging. We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section 4.2.2) to guide our parameter substitution strategy search.
For a model trained on unverified data, access to fully clean data is often not guaranteed. However, the corresponding task is typically known to the defender, providing an opportunity to identify homologous models trained on overlapping tasks that can be used for merging. For instance, previous works, WAG and PURE, utilized the IMDB (Maas et al., 2011) dataset to purify models trained on poisoned SST-2 datasets, as both share the sentiment classification task domain.
In this work, we select proxy models that share a similar domain with the poisoned tasks. Specifically, we use Twitter Abusive (Founta et al., 2018) for OLID, SNLI [^3] Creative Commons Attribution-ShareAlike 4.0 International License (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews. To evaluate generalizability, we use three sentiment classification datasets—IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)—to purify victim SST-2 models through model merging. All datasets are downloaded from Hugging Face, which adheres to the Apache License 2.0.
Our method involves only one hyperparameter, $\alpha$, as described in  Equation 1, which controls the preference between model utility and attack resistance during the model merging strategy search. We set $\alpha$ to 0.4 by default, which slightly favors seeking a more attack-resistant model. It can be adjusted to smaller values (e.g., 0.1) when a lower ASR is the primary target, and to larger values (e.g., 1.0) when high utility is demanded.
A.2 Computational Resources
We conduct experiments using three seeds on Nvidia V100 and A1000 GPUs, and report the average scores. Running our method takes only 4 minutes on a single GPU.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
ONION 56.8 92.9 99.9 93.3 85.7 91.9 92.9 92.9
Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
PURE 0 50.9 0 50.9 0 50.9 0 50.9
TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
DARE w/ TIES 99.3 96.0 100.0 96.2 96.4 95.7 92.7 95.8
WAG 84.4 94.8 60.1 95.2 58.8 94.8 56.2 92.5
SST-2 Ours (GMS) 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
Benign 29.0 84.9 31.0 84.9 45.6 84.9 53.6 84.9
Victim 99.9 85.2 100.0 85.0 94.5 85.1 100.0 85.0
Proxy Model (Twitter) 37.8 84.4 40.0 84.4 55.4 84.4 67.1 84.4
ONION 75.0 84.8 99.4 84.8 86.1 84.1 99.6 84.7
Z-Def. 29.7 85.6 30.3 85.3 93.5 85.3 53.9 85.7
PURE 82.2 75.2 87.5 75.2 76.1 79.1 100.0 72.1
TIES 36.2 84.8 38.2 84.8 57.5 84.5 65.3 84.5
DARE w/ TIES 95.6 86.3 77.8 86.0 90.5 85.9 86.8 86.3
WAG 52.0 85.0 48.6 84.6 63.1 84.5 68.2 85.0
OLID Ours (GMS) 32.1 85.0 28.0 84.2 52.2 84.4 64.2 84.9
Benign 12.3 87.6 12.6 87.6 26.4 87.6 36.9 87.6
Victim 100.0 89.4 100.0 90.3 96.0 89.0 99.9 89.4
Proxy Model (SNLI) 12.2 84.1 9.2 84.1 25.3 84.1 31.7 84.1
ONION 64.3 86.1 98.6 86.9 89.0 85.5 98.8 86.6
Z-Def. 11.1 88.3 11.6 89.7 92.2 89.1 50.6 89.7
PURE 33.3 33.8 33.3 33.8 33.3 33.8 33.3 33.8
TIES 91.6 89.2 94.5 89.5 80.7 89.2 88.6 90.0
DARE w/ TIES 93.5 90.5 100.0 91.4 93.9 89.9 99.4 90.5
WAG 71.0 88.7 60.3 88.5 77.9 88.0 80.5 88.8
MNLI Ours (GMS) 10.8 86.5 10.7 86.3 14.0 86.5 31.7 86.3
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
ONION 59.4 94.8 97.8 95.1 84.8 94.5 99.6 94.7
Z-Def. 1.6 95.3 0.4 95.3 97.9 96.1 100.0 95.0
PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
DARE w/ TIES 99.9 95.2 99.6 95.4 97.8 96.5 100.0 95.2
WAG 92.7 94.1 97.8 94.4 78.0 93.9 90.9 94.3
AGNews Ours (GMS) 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
Table 8: Performance of our method compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds.
A.3 Implementation details for baselines
We follow the open-source implementations for each baseline, and basically using their default hyper-parameters, while maintaining using the identical datasets, backdoor settings and the trained models for a fair comparison.
For the baseline DARE Yu et al. (2024), it first applies random parameter dropping and rescaling to the involved models with a specified drop rate, and then incorporates model merging techniques to combine the processed models. Various model merging methods can be integrated with DARE, and we choose TIES merging as a representative, as it demonstrates decent performance for encoder-based models (e.g., bert-base, roberta-base). Their method involves three tunable hyperparameters for encoder-based LMs, as outlined in Table 5 of their paper: drop rate, scaling term, and ratio to retain. We retained the original search space for drop rate and scaling term but expanded the ratio to retain from [0.1, 0.2, 0.3] to [0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8] for the roberta-large model to ensure more complete search.
For the baseline PURE (Zhao et al., 2024), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts. The pruned model is then fine-tuned further on the proxy clean dataset to normalize the remaining attention heads for purification. To ensure a fair comparison in the main experiments with the roberta-large model, we use the same poisoned and proxy datasets, e.g., poisoned on the SST-2 dataset with IMDB as the proxy clean dataset. We set the hyperparameter accuracy threshold (used to stop head pruning) to 0.90 for SST-2, MNLI, and AGNews, and 0.85 for OLID to prevent overly aggressive pruning. For the bert-base model, we follow the original implementation, including the use of the SST-2 dataset for fine-tuning SST-2 victim models and maintaining the default accuracy threshold of 0.85. While PURE uses the label flip rate (LFR) as its evaluation metric for backdoor defense (implanting triggers into test data while keeping the labels unchanged), we adopt the attack success rate (ASR) on label-flipped test data as our evaluation metric for a fair comparison with our method.
We adhere to the default implementations and hyperparameter settings for all other baseline methods.
Appendix B Additional results
BadNet InsertSent LWS HiddenKiller
[…]
Benign 8.5 92.9 3.5 92.9 22.2 92.9 17.3 92.9
Victim 100.0 92.9 100.0 92.2 98.1 91.5 95.9 91.8
Proxy Model (IMDB) 10.4 85.4 6.4 85.4 17.9 85.4 14.7 85.4
ONION 58.0 89.9 99.7 89.9 85.4 88.6 94.4 89.2
Z-Def. 8.3 92.5 1.8 91.9 97.4 90.9 38.7 91.5
PURE 30.1 92.0 10.7 91.5 66.4 91.1 28.7 91.3
TIES 99.2 92.7 98.2 92.4 91.2 92.3 86.8 92.6
DARE w/ TIES 100.0 93.0 90.8 88.9 92.6 92.7 94.1 90.6
WAG 74.3 91.9 70.7 91.7 73.5 91.7 60.3 91.6
SST-2 Ours (GMS) 10.8 86.8 9.8 88.8 27.4 88.0 19.5 86.8
Table 9: Performance of our method compared to baselines under various backdoor attacks on the BERT-base model, with each value averaged over three seeds. The victim model is trained on SST-2 dataset.
B.1 Performance across different model architectures and different dataset
Performance across different dataset. Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table 2. The settings are consistent with those in Section 5.2, with each score averaged over three runs using different seeds. For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure 4. The top two ASR performances are highlighted. Across all datasets and backdoor attacks, our method consistently ranks among the top two in backdoor removal performance, with minimal harm to clean accuracy. Notably, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS achieves significant improvements (at least 25%) over all baselines. For instance, on SST-2, GMS reduces the ASR for LWS to 9.7%, compared to 58.8% for the next-best baseline, Z-Def.
While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large. This instability seems to arise from PURE’s reliance on accuracy from the clean proxy dataset as the pruning stopping criterion. For well-trained models with high accuracy, a substantial number of heads are pruned, leading to a broken purified model (e.g., 0% ASR but random-guess-level CACC).
Performance across different model architectures. We also evaluated our method on different architectures. Table 9 shows the results on the BERT-base model. The poison rate remains at 20%, and each model is trained for three epochs. Scores are reported under $\alpha=0.4$ and averaged over three seeds: 1000, 2000, and 3000. Across all attacks, our method performs consistently well, particularly excelling against more challenging attack strategies. Notably, it achieves over a 39% advantage in defending against LWS and a 9% improvement over the second-best baseline for HiddenKiller.
B.2 Performance under different poison rate: 20%, 10% and 5%
While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table 10. Our method is effective in model purification across all poison rates.
Victim Proxy Poison Rate Metric Result
[…]
Table 10: Results for SST-2 with varying poison rates.
Figure 3: Results of using different weights (alpha) on OLID dataset.
Figure 4: Results of using different weights (alpha) on AGNEWS dataset.
B.3 How sensitive is our method to the selection of hyperparameters?
Our method involves two hyperparameters: $\alpha$ in Equation 1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results. Below, we justify these choices. As shown in Figure 4 and Figure 4, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength. An exception is the HiddenKiller attack on the OLID dataset, where a more aggressive $\alpha=0.1$ is recommended. This is because OLID is a dataset collected from tweets, while HiddenKiller paraphrases the data using formal syntactic templates, drastically altering OLID’s language style. This change significantly impacts CACC, necessitating a higher weight for backdoor removal to ensure effective purification. For stopping patience, as shown in Figure 5, the score computed in Equation 1 generally follows a (weakly) monotonically decreasing trend. While a larger $T$ improves performance, we choose $T=5$ as a reasonable balance between performance and efficiency.
Figure 5: The search iteration history for four kinds of backdoors on SST-2 dataset.
Methods
Metrics Stage WAG PURE Clean Ours
Before 94.72 93.12 95.76 91.86
[…]
Before 84.40   6.31   4.05   3.83
[…]
Table 11: Evaluating the robustness of different parameter purification methods towards the Retuning Attack (RA) using 0.1% of the original poisoned training set (which contains 20% poisoned data) for fine-tuning. We report the performance of each methods before and after the RA.
Figure 6: The optimal substitution strategy for defending against each backdoor attack for the roberta-large model trained on the SST-2 dataset. The green squares indicate the substituted modules.
B.4 Robustness to Retuning Attacks
Recent works Min et al. (2024); Qi et al. (2024) have highlighted a critical pitfall in backdoor defenses: purified models with low ASR often fail to completely eliminate inserted backdoor features. A straightforward Retuning Attack (RA)—which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)—can easily recover the backdoor in the purified model, as demonstrated in Table 11.
Several related studies Qi et al. (2024); Chen et al. (2024b); Tarun et al. (2024) have sought to uncover the underlying reasons for this failure. A key observation is that the purified model’s parameters do not sufficiently deviate from those of the backdoored model. Specifically, the weight differences between a purified model and a (retrained) benign model are significantly larger than those between the purified and backdoored models. This suggests that purified and backdoored models are connected via a backdoor-related path in the loss landscape Min et al. (2024). Such vulnerabilities pose severe risks, as purified models are often deployed in various downstream applications, and even with backdoor defenses applied, attackers can easily re-trigger the backdoor in downstream tasks Min et al. (2024).
This challenge motivates us to propose a purification method aimed at breaking this backdoor-connected path. Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024) or merging them with other parameters Arora et al. (2024); Yadav et al. (2023). Moreover, because parameters often interact within functional modules, modifying at the module level rather than the individual parameter level tends to remove backdoor features more effectively. As shown in Table 11, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al. (2024), the parameter-merging baseline, WAG Arora et al. (2024), is easily re-triggered. In contrast, our module-substitution method demonstrates greater robustness while maintaining utility, performing comparably to the clean model baseline (i.e., finetuning a clean model on the retuning dataset). We provide our baseline PURE with additional advantages by granting access to the full clean dataset of the victim task for finetuning, as obtaining a usable purified model otherwise proves challenging. While the recent method PURE also demonstrates competitive performance, it is worth noting that significant effort was required to find a configuration where its pruning step did not break the purified model. Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%—nearly twice the robustness.
B.5 Examples of Searched Substitution Strategy
We applied our GMS defense to purify victim models compromised by various backdoor attacks across four datasets. Examples of the searched strategies from the SST-2 dataset are shown in Figure 6. While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section 5.3 and  Section B.6.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
Benign 29.0 84.9 31.0 84.9 45.6 84.9 53.6 84.9
Victim 99.9 85.2 100.0 85.0 94.5 85.1 100.0 85.0
Proxy Model (Twitter) 37.8 84.4 40.0 84.4 55.4 84.4 67.1 84.4
GMS 32.1 85.0 28.0 84.2 52.2 84.4 64.2 84.9
OLID OLID BadNet Strategy 33.6 85.0 37.4 84.6 56.5 84.3 65.0 84.7
Benign 12.3 87.6 12.6 87.6 26.4 87.6 36.9 87.6
Victim 100.0 89.4 100.0 90.3 96.0 89.0 99.9 89.4
Proxy Model (SNLI) 12.2 84.1 9.2 84.1 25.3 84.1 31.7 84.1
GMS 10.8 86.5 10.7 86.3 14.0 86.5 31.7 86.3
MNLI MNLI BadNet Strategy 12.2 86.3 10.8 86.3 27.0 85.2 35.2 86.4
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
Table 12: The performance of transferring strategy searched based on BadNet to other attacks.
B.6 Transferability of substitution strategies
We present the results for SST-2 and AGNews in Table 7, with additional results for other datasets provided in Table 12. All experiments follow a consistent setup: the substitution strategy is optimized based on BadNet attacks and then applied to coordinate module substitutions for other attacks. For instance, in the MNLI task, the substitution strategy identified for defending against BadNet—using a proxy model trained on SNLI—is directly applied to replace modules in InsertSent, LWS, and HiddenKiller with those from the proxy model. Across all tasks, the defense performance is surprisingly strong, revealing an intriguing observation: a substantial overlap exists in the components impacted by different attacks.

### v2

@@
-Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution
+Cut the Deadwood Out: Backdoor Purification via Guided Module Substitution
 Abstract
-The success of DNNs often depends on training with large-scale datasets, but building such datasets is both expensive and challenging.
-Consequently, public datasets from open-source platforms like HuggingFace have become popular, posing significant risks of data poisoning attacks.
-Existing backdoor defenses in NLP primarily focus on identifying and removing poisoned samples; however, purifying a backdoored model with these sample-cleaning approaches typically requires expensive retraining.
-Therefore, we propose Greedy Module-Substitution (GMS), which identifies and substitutes “deadwood” modules (i.e., components critical to backdoor pathways) in a backdoored model to purify it.
-Our method relaxes the common dependency of prior model purification methods on clean datasets or clean auxiliary models.
-When applied to RoBERTa-large under backdoor attacks, GMS demonstrates strong effectiveness across various settings, particularly against widely recognized challenging attacks like LWS, achieving a post-purification attack success rate (ASR) of 9.7% on SST-2 compared to 58.8% for the best baseline approach.
-Cut the Deadwood Out: Post-Training Model Purification
-with Selective Module Substitution
+Model NLP models are commonly trained (or fine-tuned) on datasets from untrusted platforms like HuggingFace, posing significant risks of data poisoning attacks.
+A practical yet underexplored challenge arises when such backdoors are discovered after model deployment, making retraining-required defenses less desirable due to computational costs and data constraints.
+In this work, we propose Guided Module Substitution (GMS), an effective retraining-free method based on guided merging of the victim model with just a single proxy model.
+Unlike prior ad-hoc merging defenses, GMS uses a guided trade-off signal between utility and backdoor to selectively replaces modules in the victim model.
+GMS offers four desirable properties: (1) robustness to the choice and trustworthiness of the proxy model, (2) applicability under inaccurate data knowledge, (3) stability across hyperparameters, and (4) transferability across different attacks.
+Extensive experiments on encoder models and decoder LLMs demonstrate the strong effectiveness of GMS.
+GMS significantly outperforms even the strongest defense baseline, particularly against challenging attacks like LWS.
+The code is available at https://github.com/weijun-l/guided-module-substitution.
+Cut the Deadwood Out: Backdoor Purification
+via Guided Module Substitution
 Yao Tong1thanks: Equal contribution., Weijun Li2∗, Xuanli He3, Haolan Zhan4, Qiongkai Xu2thanks: The corresponding author.
@@
 1 Introduction
-Deep neural networks (DNNs) have achieved remarkable success in various domains Krizhevsky et al.
-(2012); Bowman et al.
-(2016), largely driven by training on large-scale datasets.
-Given the high cost of collecting and labeling data from scratch, incorporating public datasets from open-source platforms like HuggingFace in training has become a common practice, particularly in NLP Howard and Ruder (2018); Ding et al.
+Modern NLP models are frequently adapted to specific downstream tasks through fine-tuning on custom datasets Howard and Ruder (2018).
+In practice, these datasets are often collected from diverse sources, some of which may be unreliable (e.g., open repositories like HuggingFace Lhoest et al.
+(2021)).
+As a result, models can unknowingly incorporate poisoned data, leading to backdoor vulnerability Zhang et al.
+(2021); Xu et al.
+(2021): a backdoored model, trained with poisoned data, behaves normally on clean inputs but misbehaves when exposed to the trigger Bai et al.
 (2025).
-However, this growing reliance on open-source datasets from unaudited platforms raises significant security concerns Yang et al.
-(2021a); Zhang et al.
-(2021), especially for commercial use.
-Malicious actors can exploit this ecosystem by publishing poisoned datasets Xu et al.
-(2021) to embed backdoors into models for potential benefit.
-These backdoors allow attackers to manipulate models, causing targeted misbehavior when triggered by specific inputs while operating normally on clean inputs.
-As a result, they often remain stealthy until their triggers are activated.
-Figure 1: Parameter-level editing (left) prunes parameters "most relevant" to backdoor features but leaves many backdoor-related parameters intact within each module, allowing backdoor pathways to potentially persist.
-In contrast, our module-level substitution (right) effectively disrupts these pathways by completely replacing backdoor-connected modules with counterparts from a proxy model (not required to be clean).
-Existing defense strategies against backdoor attacks in NLP primarily focus on backdoor samples detection, such as detecting and removing poisoned samples leveraging their distinct characteristics compared to clean ones Qi et al.
-(2021a); Yang et al.
-(2021b); Tran et al.
-(2018); He et al.
-(2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al.
-(2022, 2023); Xu et al.
-(2024).
-However, their effectiveness is often suboptimal Min et al.
-(2024), especially as backdoor triggers evolve to be stealthier and harder to detect or reverse Qi et al.
-(2021b).
-Moreover, while these methods follow a fine-tuning or adversarial training for unlearning Wei et al.
-(2023) to sanitize poisoned models, they cannot effectively purify a backdoored model without incurring a significant cost of retraining Wu et al.
-(2022).
-Then a practical question arises: How can we cost-efficiently eliminate the backdoor features from an existing model after users reported some compromise?
-While only a few model purification defenses in NLP exist, they predominantly focus on parameter-level editing.
-For instances, Fine-mixing Zhang et al.
-(2022) and Fine-purification Zhang et al.
-(2023) mix a clean pre-trained language model (PLM) with a backdoored model and rely on fine-tuning with a clean dataset for purification.
-Similarly, Zhao et al.
-(2024) prunes attention heads potentially affected by poisoned data and fine-tunes the pruned model using a clean dataset.
-However, recent studies have observed that backdoor features are encoded as complex pathways in victim models, with significant overlap between backdoor and clean neurons at the module level Jebreel et al.
+A critical yet underexplored problem arises when such backdoors are discovered after model deployment.
+Existing defenses, such as data filtering and retraining (He et al., 2024), pruning followed by fine-tuning Liu et al.
+(2018); Zhao et al.
+(2024b), or unlearning backdoors based on clean Li et al.
+(2023) or poisoned data Min et al.
 (2024); Li et al.
-(2024).
-As illustrated in Figure 1 (Left), parameter-level editing will leave hidden vulnerabilities in the non-pruned connections, thus making the defense effectiveness heavily relies on fine-tuning with a clean dataset to restore clean features Liu et al.
-(2018), an assumption that is both strong and rarely feasible in practice.
-Here, we develop a method that effectively disrupts the model’s backdoor pathways without relying on the assumption of access to a clean dataset.
-Note that the abundance of public platforms for open-source models (e.g., HuggingFace) offers semantically similar yet independent models for any task domain Semwal et al.
-(2018).
-Therefore, we propose a simple yet effective module-level purification defense, Greedy Module-Substitution (GMS), which identifies critical modules for backdoor propagation and purifies them by replacing the parameters with those of a proxy model.
-While this approach eliminates the need for a clean dataset to restore semantic knowledge, the proxy model is not required to be clean either, since the pathways of two different backdoored models are less likely to be connected (empirically verified in Table 5).
-The primary contributions of this work are as follows.
+(2021b); Chen et al.
+(2022), typically incur substantial computational costs Wu et al.
+(2022)–especially for large-scale models or proprietary pipelines.
+This raises a practical and timely question: Can we effectively purify a trained backdoored model without retraining?
+Recent work (Arora et al., 2024) has proposed model merging as a cost-efficient and retraining-free defense method for model purification: leveraging the abundance of open-source proxy models online [^1] For reference, on Hugging Face, there are 74 LLaMA-7B and 262 BERT-base models trained on SST-2 alone, and the pool of proxy models is even larger when considering all datasets within the same task domain., this approach merges several proxy models–backdoored or not–with the victim model.
+While promising, this approach relies on ad hoc merging of multiple models with no principled selection, such as utility and defense performance.
+Its effectiveness is thus unpredictable and sensitive to the number of models involved.
+Our later experiments in Section˜5.2 demonstrate that when only a single proxy model is available, such defenses become ineffective.
+More seriously, In practice, collecting more homogeneous proxy models also requires more effort and often harm utility more Zhou et al.
+(2025a); Wang et al.
+(2025b), even though more merged parameters can help neutralize backdoors.
+We therefore target a more challenging scenario–designing guided merging using only one proxy model.
+In this work, we propose a simple yet effective alternative: guided merging with a single proxy model.
+Rather than blindly merging models, we explicitly guide the merging process using a signal that approximates the trade-off between utility and backdoor risk.
+Intuitively, if retaining a particular module from the victim model contributes significantly more to the attack performance than to task utility, it is likely critical for encoding backdoor features.
+We thus progressively identify such modules and replace them with their counterparts from the proxy model, as long as the resulting utility degradation remains limited relative to backdoor mitigation gain.
+We refer to this method as Guided Module-Substitution (GMS).
+While training-free and more efficient than prior model-merging-based defenses, our method GMS not only retains the robustness of model-merging defenses to proxy model trustworthiness, but is also tolerant of imperfect data knowledge, stable across hyperparameters, and exhibits strong transferability across different attacks.
+Further discussion of these desirable properties is in Section˜5.3.
+The contributions of this work are as follows.
   1. 1.
-    We propose the first efficient (quadratic time complexity) module-level backdoor purification defense (GMS), which relaxes the strong reliance on clean auxiliary models and clean datasets required by prior defenses, as shown in Table 5 and Table 6.
+    We propose GMS, the first backdoor purification method based on guided merging of a victim model with a single proxy model.
+This design enables purification without retraining while remaining efficient and practical.
   2. 2.
-    Our method significantly outperforms baseline approaches against challenging backdoor attacks like LWS and HiddenKiller, e.g., on SST-2 victim model, GMS reduces the ASR on LWS to 9.7%, compared to 12.8% for the benign model and 58.8% for the best baseline.
+    Experiments on standard NLP models and LLMs (Section˜5.2) show that GMS consistently surpasses competitive defense baselines, with especially strong performance against the more challenging attacks like LWS and HiddenKiller.
   3. 3.
-    Our GMS purification strategy exhibits strong transferability across different attacks.
-Even the strategy with few module substitutions can effectively defend against other backdoor attacks, making it suitable when the victim dataset is inaccessible (Section 5.3 and Table 7).
+    We further validate four desirable properties of GMS in Section˜5.3 and Sections˜B.4, B.5, B.6, B.7, B.8, B.9, B.10, B.11 and B.12, namely (1) robustness to proxy model choices, (2) applicability to varied data conditions, (3) stability across hyperparameters, and (4) transferability across attacks.
 2 Related works
-Backdoor attack
+Backdoor attack.
 Backdoor attacks aim to manipulate models to behave normally on benign inputs while exhibiting attacker-controlled behavior when specific triggers are present.
@@
 (2019); Qi et al.
-(2021d, c) with triggers range from misspelled words Chen et al.
+(2021c, b) with triggers range from misspelled words Chen et al.
 (2021), rare words Yang et al.
@@
 (2020) and syntactic structures Qi et al.
-(2021c) to text styles Pan et al.
+(2021b) to text styles Pan et al.
 (2022), posing serious challenges for detection.
@@
 In this work, we focus on data-poisoning backdoor attacks in NLP due to their widespread applicability and practical implications.
-Backdoor defense
-Backdoor defense approaches can be categorized according to their strategies: (1) Backdoor sample detection focuses on identifying poisoned samples in the dataset.
-The defense approaches usually include detecting and removing backdoor samples, leveraging the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones Qi et al.
+Backdoor defense.
+Post-training backdoor defense approaches can be broadly categorized as: (1) Backdoor sample detection followed by training, which first filters poisoned samples from the dataset and then retrains or fine-tunes the model.
+These defenses typically rely on the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones, enabling their detection Qi et al.
 (2021a); Yang et al.
@@
 (2018); He et al.
-(2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al.
+(2023), or attempt to reverse-engineer backdoor triggers to neutralize their impact Wang et al.
 (2019, 2022, 2023); Xu et al.
 (2024).
-These strategies are typically applied during the training phase or involve fine-tuning after removing backdoor samples.
-(2) Backdoor model purification: These methods aim to eliminate backdoor features from a well-trained model.
+Afterward, such strategies usually involve unlearning the identified backdoor samples or retraining/fine-tuning on the filtered dataset.
+However, their effectiveness is often unsatisfactory Sun et al.
+(2025); Wang et al.
+(2023); Qi et al.
+(2021a), primarily because accurately identifying all poisoned samples is (increasing) difficult Wu et al.
+(2022); Qi et al.
+(2021a); Yang et al.
+(2021b); Tran et al.
+(2018); He et al.
+(2023), leaving residual backdoor behavior even after retraining or unlearning.
+Our method, in this regard, compensates for these shortcomings by being more robust to data filtering quality, and thus provides a more reliable alternative to unlearning- or retraining-based defenses.
+(2) Backdoor model purification: These methods aim to eliminate backdoor features directly from a well-trained model and are generally regarded as achieving state-of-the-art defense performance Zhu et al.
+(2023); Zhao et al.
+(2024b).
 Common approaches involve merging parameters with other proxy models Arora et al.
 (2024); Chen et al.
-(2024a), or fine-tuning or pruning using a limited amount of clean data Wu and Wang (2021); Min et al.
+(2024a), or pruning and fine-tuning using a limited amount of clean data Wu and Wang (2021); Min et al.
 (2023); Liu et al.
 (2018); Zhao et al.
-(2024); Zhu et al.
+(2024b); Zhu et al.
+(2023), which may sometimes be assisted by clean proxy models as well Liu et al.
+(2018); Zhang et al.
 (2023).
-These techniques not only reduce training costs but also achieve state-of-the-art defense performance Zhu et al.
-(2023); Zhao et al.
-(2024); Arora et al.
+Our work advances the second line of research by minimizing the number of auxiliary models to one and by eliminating the need for additional training and strictly clean datasets.
+Model merge.
+Model merging methods have emerged as an efficient way to build powerful models by combining existing trained domain-specialized models without requiring extensive retraining Wortsman et al.
+(2022); Ilharco et al.
+(2023); Cheng et al.
 (2025).
-Our work advances the second line of research by minimizing the number of auxiliary models to one and reducing dependence on the clean dataset.
+Most advances focus on developing utility-preserving multi-task models (e.g., merging models trained on different tasks like coding and mathematics He et al.
+(2025)), or stronger domain-specific models by combining within-domain models Wortsman et al.
+(2022).
+Surprisingly, recent work shows that naive weight averaging across multiple models can mitigate backdoors Arora et al.
+(2024).
+Yet such blind merging is inefficient–requiring unpredictable trial-and-error in model collection, since one cannot know in advance how many models are needed for purification–and often harmful, as merging many models can yield utility drops from parameter interference and redundancy Zhou et al.
+(2025a); Wang et al.
+(2025b).
+Thus, our work address these issues by proposing effective and efficient guided merging with a single proxy model.
+Note that while there are some works studying how model merging systems themselves can be exploited for backdooring, leveraging the fact that merging methods aim to preserve (task) parameters from all models Wang et al.
+(2025a); Zhang et al.
+(2024); Yuan et al.
+(2025), our objective is fundamentally different: we use merging as a tool to purify a single target model, where only the target model’s clean task is relevant and other tasks can all be discarded.
+Since our main goal is to demonstrate guided merging can act as a defense–and obtaining a single homogeneous proxy is practical–we focus on the standard homogeneous setting.
+As heterogeneous merging itself remains an open challenge Xu et al.
+(2025); Du et al.
+(2025), exploring its use for defense is left to future work.
+Safety Localization.
+Recent studies suggest that safety-critical behaviors in LLMs are often localized to specific layers or components.
+For example, jailbreak defenses identify safety layers that are critical for aligning harmful queries Zhao et al.
+(2024a); Ouyang et al.
+(2025); Zhou et al.
+(2024); a small set of safety layers Li et al.
+(2025) or safety modules (e.g., attention heads) Zhou et al.
+(2025b) are crucial for distinguishing malicious from benign inputs; and modifying specific safety-related parameters can further enhance alignment Wei et al.
+(2024).
+Similarly, backdoor research shows that certain layers Jebreel et al.
+(2023) and even neurons or heads Zhao et al.
+(2024b) disproportionately influence attack success, enabling defenses through targeted editing or masking.
+These findings motivate our focus on module- and layer-level guided merging, which balances efficiency with effectiveness.
 3 Preliminary
@@
   1. 1.
-    Layer Set ($\mathcal{L}$) is the set of block indices, capturing the architecture’s depth and structure.
+    Layer Set ($\mathcal{L}$) is the set of block indices that captures the depth of the model.
   2. 2.
     Module Set ($\mathcal{M}$) is the set of representative functional modules within each block.
-Transformer blocks
+Transformer blocks.
 In this paper, we use Transformer-based architecture Vaswani et al. (2017) as the testbed, given its widespread attention and implementation in NLP.
 For each Transformer block, the input will be processed through Attention and Feed-Forward Network (FFN) modules.
-Attention modules contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively.
+An attention module contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively.
 The following FFN contains two components: $W_{F}$ for the forward layer and $W_{P}$ for the projection layer, connected by a non-linear activation function.
 In total, each Transformer block contains six key functional modules with parameters $W_{Q},W_{K},W_{V},W_{O},W_{F},$ and $W_{P}$.
-For simplicity and consistency, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper.
-Accordingly, we define the module set $\mathcal{M}$ for Transformer models as $\{Q,K,V,O,F,P\}$ and layer set =$\mathcal{L}=\{1,2,\ldots,L\}$.
-Evaluation metrics
-Following previous literature (Qi et al., 2020), we adopt two metrics: (1) Clean Accuracy (CACC) measures the prediction accuracy on clean samples, and (2) Attack Success Rate (ASR) measures the attack accuracy on poisoned test set.
-A higher CACC reflects better utility, while lower ASR indicates stronger defense.
+For simplicity, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper.
+Accordingly, we define the module set for Transformer models $\mathcal{M}=\{Q,K,V,O,F,P\}$ and layer set $\mathcal{L}=\{1,2,\ldots,L\}$.
 Figure 1: The pipeline of our method.
-Step 1: Extract two small proxy datasets from the suspected dataset $\mathcal{D}$ used for computing score in Equation 1.
+Step 1: Extract two small proxy datasets used for computing score in Equation˜1.
 Step 2: Iteratively update a substitution matrix $S$ to greedily maximize the trade-off score between backdoor removal and utility preservation.
 Step 3: Return purified model $M_{pure}$ corresponding to the best substitution matrix $S_{best}$.
-For more details, refer to Algorithm 1.
+For more details, refer to Algorithm˜1.
 4 Methodology
-In Section 4.1, we introduce the problem setting, and in Section 4.2 we present our methods.
+In Section˜4.1, we introduce the problem setting, and we present our methods in Section˜4.2.
 4.1 Problem setting
-Threat Model
-We consider the scenario of data poisoning-based backdoor attacks, where an attacker releases a poisoned dataset $\mathcal{D}$ to embed backdoors into any model trained on this dataset.
-Starting with a clean dataset $\mathcal{D}_{\text{cl}}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ denote the input and label spaces, respectively, the attacker selects a subset $\mathcal{D}_{\text{sub}}\subset\mathcal{D}_{\text{cl}}$, and then modified each sample $(x,y)\in\mathcal{D}_{\text{sub}}$ into a poisoned sample $(\tilde{x},y_{t})$ where $\tilde{x}=g(x,\Delta)$, the pre-defined trigger is $\Delta$, $g$ is a generation function, and $y_{t}$ is the target label.
-The poisoned subset is $\mathcal{D}_{\text{poi}}=\{(\tilde{x},y_{t})\mid(x,y)\in\mathcal{D}_{\text{sub}}\}$.
-Finally, the attackers publish the poisoned dataset constructed as $\mathcal{D}=(\mathcal{D}_{\text{cl}}\setminus\mathcal{D}_{\text{sub}})\cup\mathcal{D}_{\text{poi}}$.
-Defense Scenario
-The defender (e.g., a model vendor) has already trained a (potentially backdoored) victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$.
-While the defender has access to $\mathcal{D}$, they are unaware of all possible backdoor patterns.
-[^1] As shown in Section 4.2.2, the full training dataset is not strictly required; a small subset of poisoned data reported by users or automatically detected by algorithms is sufficient for applying our defense.
-The defender has white-box access to a trained victim model $M_{v}$ and a proxy model $M_{p}$ trained on similar downstream tasks.
+Defense Setting.
+The defender (e.g., a model vendor) has trained a potentially backdoored victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$, which may contain poisoned samples.
+Following the standard setting Li et al.
+(2021b), we assume the defender has full control over the training process but lacks prior knowledge of the backdoor patterns, their proportion, or their distribution within $\mathcal{D}$.
+If an isolation method is employed, it may only identify a subset of the poisoned examples, as perfect detection is not guaranteed.
+The defender has white-box access to the trained victim model $M_{v}$, as well as access to a proxy model $M_{p}$ trained on similar downstream tasks.
 4.2 Our method
-We aim to disrupt the backdoor connection in victim model $M_{v}$ while preserving the knowledge required for the original task.
-Models trained on similar tasks tend to capture similar domain-specific knowledge Semwal et al.
-(2018).
-Hence, we replace backdoor-connected modules in the $M_{v}$ with counterparts from a proxy model $M_{p}$ that shares similar domain knowledge.
-Notably, $M_{p}$ does NOT need to be benign, as even a malicious proxy model lacking identical backdoor pathways will fail to restore the disrupted backdoor connection.
-Our main algorithm, Greedy Module-Substitution (GMS), is detailed in Section 4.2.3.
-Section 4.2.1 defines the objective optimized by GMS, and Section 4.2.2 describes the method to construct the proxy datasets used in evaluation.
+We aim to combine a victim model with a proxy model.
+Unlike traditional model merging Arora et al.
+(2024), which prioritizes utility across all models’ tasks, our method targets backdoor removal and requires preserving only the clean task performance of the target model.
+This enables a more targeted substitution strategy: directly replacing backdoor-encoded modules in the victim model $M_{v}$ with counterparts from the proxy model $M_{p}$.
+Specifically, we operate at the module- and layer-level, inspired by recent findings that malicious behaviors in LLMs are often localized to specific safety layers and safety modules (discussed in Section˜2).
+While not the only possible granularity, this level strikes a balance between efficiency and effectiveness, and offers greater robustness and transferability than neuron-level edits, as discussed in Section˜4.2.3 and Section˜5.3.
+Our main algorithm, Guided Module-Substitution (GMS), is detailed in Section˜4.2.3.
+Section˜4.2.1 defines the objective optimized by GMS, and Section˜4.2.2 provides an example of constructing proxy datasets.
 4.2.1 Objective
-The defender’s objective is to identify a purified model $M_{pure}$ from all possible substituted models $M$ that effectively eliminates the backdoor while preserving task utility.
-This can be formulated as
+The defender’s objective is to identify a purified model $M$ within the model space $\mathbb{M}$, defined by all possible substitution strategies over the layer set $\mathcal{L}$ and module set $\mathcal{M}$, that effectively eliminates the backdoor while preserving task utility:
 $\operatorname*{argmax}_{M}\,(1-\alpha)\cdot\Delta_{\text{asr}}(M)+\alpha\cdot(1-\Delta_{\text{acc}}(M))$ (1)
-where $\delta_{\text{asr}}(M)$ measures the extent of backdoor removal by quantifying the reduction in ASR and $\delta_{\text{acc}}(M)$ assesses the impact on task utility by measuring the decrease in CACC.
-However, since the defender lacks access to the clean dataset and poisoned test set, we adopt the method described in Section 4.2.2 to extract two proxy datasets, $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, from the suspected dataset $\mathcal{D}$ for ASR and CACC estimation.
-The goal is to maximize $\delta_{\text{asr}}(M)$ while minimizing $\delta_{\text{acc}}(M)$.
-The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor defense and task utility.
+where $\Delta_{\text{asr}}(M)$is a score that reflects the extent of backdoor removal and $\Delta_{\text{acc}}(M)$ assesses the reduction in task utility.
+Specifically, if given access to clean and poisoned samples, we can evaluate the attack success rate (ASR) on poisoned data and the clean accuracy (ACC) on benign inputs.
+The improvement in backdoor defense and degradation in utility for a given model $M$ relative to the original victim model $M_{v}$ are defined as:
+$\Delta_{\text{asr}}(M)$ $=s_{asr}(M_{v})-s_{asr}(M),$ (2)
+$\Delta_{\text{acc}}(M)$ $=s_{acc}(M_{v})-s_{acc}(M).$ (3)
+The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor removal and task utility preservation.
 A larger $\alpha$ favors utility, while a smaller $\alpha$ prioritizes backdoor defense.
 4.2.2 Proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$
-In this section, we introduce the process of extracting two proxy datasets from the suspected dataset $\mathcal{D}$: a proxy clean set $\mathcal{D}_{clean}$, which is identified with high probability as clean (i.e., they are likely to be clean though not guaranteed to be free from poisoning), and a proxy poison set $\mathcal{D}_{poison}$, which is identified with high confidence as being poisoned.
-Both proxy datasets are not required to be purely clean or poisoned.
-When the defender has no prior knowledge of the poisoned samples, various methods can be used to construct proxy datasets, such as sample diagnosis Gao et al.
+To approximate Equations˜2 and 3, we extract two proxy datasets: a proxy clean set $\mathcal{D}_{\text{clean}}$ and a proxy poison set $\mathcal{D}_{\text{poison}}$, without assuming any prior knowledge of the backdoor.
+We only follow the standard setting to assume access to the (potentially poisoned) training set $\mathcal{D}$, or to a significantly smaller subset (e.g., 1%) of it (e.g., in user-reporting scenarios).
+In literature, various methods can be used to construct proxy datasets under such settings, such as sample diagnosis Gao et al.
 (2019); Huang et al.
@@
 (2024).
-If additional prior knowledge, such as a user-reported clean dataset, is available (as tested in Table 6), the performance of our method is expected to be improved.
-For simplicity, in this paper, we use random sampling to construct the proxy clean set $\mathcal{D}_{clean}$, based on the practical assumption that the poison samples will not dominate a training set.
-To construct the proxy poison set $\mathcal{D}_{poison}$, we adopt an existing backdoor sample detection method from SEEP He et al.
-(2024), which leverages an observation that backdoored models tend to exhibit over-confidence on poisoned samples Li et al.
-(2021b).
-SEEP utilizes training dynamics to identify anomalous samples in the training corpus, which are proven to have a high probability of being backdoor samples.
-In our example, we only focus on the most confident samples extracted (i.e., seed set).
-In Section 5.2, we show that extracting only 1% of the original poisoned data in $\mathcal{D}_{poison}$ and allowing even 20% poisoned data in $\mathcal{D}_{clean}$ are sufficient to launch an effective defense using our Greedy-Module-Substitution (GMS).
+For simplicity, we adopt the following heuristics in this paper: we construct $\mathcal{D}_{\text{clean}}$ via random sampling, assuming that poisoned samples do not dominate the dataset; and we construct $\mathcal{D}_{\text{poison}}$ using a naive confidence-based heuristic, based on the observation that poisoned examples often yield higher output confidence scores on backdoored models Li et al.
+(2021b); Swayamdipta et al.
+(2020).
 4.2.3 Greedy search for purified model $M_{pure}$
-Using proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$, we can estimate ASR and ACC to quantify the level of backdoor defense and task utility for a model $M$, we define the following two metrics:
-$\delta_{\text{asr}}(M)$ $=s_{asr}(M_{v})-s_{asr}(M),$ (2)
-$\delta_{\text{acc}}(M)$ $=s_{acc}(M_{v})-s_{acc}(M).$ (3)
-Here, $s_{acc}(M)$ represents the CACC of $M$ on $\mathcal{D}_{clean}$, while $s_{asr}(M)$ denotes its ASR on $\mathcal{D}_{poison}$.
-Similarly, $s_{acc}(M_{v})$ and $s_{asr}(M_{v})$ are the ASR and CACC for the original victim model $M_{v}$.
-In this context, $\delta_{\text{asr}}(M)$ quantifies the reduction in ASR achieved by a purified model $M$, while $\delta_{\text{acc}}(M)$ reflects the drop in task utility.
-Plugging them back into the objective (Equation 1) gives us an score.
-[^2] An inaccurate $\mathcal{D}_{poison}$ may lead to a smaller $\delta_{\text{asr}}$, while an impure $\mathcal{D}_{clean}$ might result in a larger $\delta_{\text{acc}}(M)$.
-However, since the metrics are always calculated on fixed proxy datasets, these effects can be controlled by adjusting $\alpha$.
-While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\cdot 12}$ for a BERT-base (Devlin et al., 2019)).
-Thus, we propose a greedy algorithm that transforms the problem into a feasible search by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
-Algorithm 1 Greedy-Module-Substitution (GMS)
+Using the extracted proxy datasets $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, we compute (potentially inaccurate) estimates of Equations˜2 and 3222An imperfect $\mathcal{D}_{\text{poison}}$ will lead to an underestimated $\Delta_{\text{asr}}(M)$, while an impure $\mathcal{D}_{\text{clean}}$ results in an overestimated $\Delta_{\text{acc}}(M)$.
+However, what matters is only the relative scaling between these two after weighting by $\alpha$ (Appendix C).and plug into the objective Equation˜1 to get the scoring function that guides the substitution process.
+While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\times 12}$ for a BERT-base model (Devlin et al., 2019)).
+Moreover, considering individual parameters as functional units for editing often falls short in breaking backdoor connections (thereby leading prior purification works to heavily rely on clean fine-tuning to restore clean features Liu et al.
+(2018)), we propose a greedy algorithm that transforms the problem into a feasible search–balancing granularity and scalability by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
+Algorithm 1 Guided-Module-Substitution (GMS)
 1:Input: Victim model $M_{v}$, proxy model $M_{p}$, module set $\mathcal{M}$, layer set $\mathcal{L}$, proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
@@
 3: Initialize $S\leftarrow\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$, $c_{best}\leftarrow$-$\infty$, $S_{best}\leftarrow\varnothing$
-4:  $(s_{acc},s_{asr})\leftarrow$ Evaluate($M_{v}$)
+[…]
 5:# Iteratively Update Selected Modules and Layers:
@@
 24:return $M_{pure}=\text{Replace}(M_{v},M_{p},S_{best})$
-Purifying $M_{v}$ by module substitution
-Given a backdoored model $M_{v}$ (characterized by both high CACC on a clean dataset and high ASR on a poisoned dataset), we aim to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\delta_{\text{asr}}$).
-Our algorithm, as demonstrated in Figure 2 and Algorithm 1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if =$\mathcal{M}=\{O,F,P\}$ and =$\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$.
+Purifying $M_{v}$ by module substitution.
+Given a victim model $M_{v}$ (likely to has high ACC and high ASR), we next illustrate how to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\Delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\Delta_{\text{asr}}$).
+Our algorithm, as demonstrated in Figure˜1 and Algorithm˜1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if $\mathcal{M}=\{O,F,P\}$ and $\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$.
 For simplicity, we introduce a substitution matrix $S\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{L}|}$ to specify which modules in which layer in the victim model $M_{v}$ are selected for substitution.
@@
 Given $S$, we can succinctly represent the parameter substitution as $S\odot M_{p}+(1-S)\odot M_{v}$, where $\odot$ denotes the element-wise product.
-We start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., =$\mathcal{M}=\{K,Q,V,O,F,P\}$, =$\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$.
+Since the primary goal is backdoor removal, we start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., $\mathcal{M}=\{K,Q,V,O,F,P\}$, $\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$.
 At $t$-th iteration, the algorithm considers two types of updates to $S^{t}$: deselecting a module from the current module set $\mathcal{M}$ (affecting all selected layers) or deselecting a layer from the layer set $\mathcal{L}$ (affecting all selected modules).
-The update that maximizes the score in Equation 1 is applied to $S^{t+1}$.
+The update that maximizes the score in Equation˜1 is applied to $S^{t+1}$.
 The process continues until one of two conditions is met: (1) further iterations will not yield a better strategy (controlled by stopping patience $T$); or (2) no additional layers or modules can be removed, i.e., $|\mathcal{M}|=|\mathcal{L}|=1$.
@@
 These datasets cover text classification and natural language inference (NLI) tasks, including binary and multi-class classification scenarios, and are widely used for evaluating text backdoor attacks and defenses (Qi et al., 2021a; He et al., 2023; Gupta and Krishna, 2023; Arora et al., 2024).
-We download the open-source versions of these datasets provided by HuggingFace (Lhoest et al., 2021).
-Table 1 presents the statistics of the evaluated datasets.
-To train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively.
-For MNLI, we use a random subset of 100,000 samples to reduce overhead.
+We adapt the open-source datasets provided by HuggingFace (Lhoest et al., 2021).
+Table 1 are the statistics of these datasets.
 Backdoor Attacks.
-We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021d), and (4) HiddenKiller (Qi et al., 2021c).
+We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021c), and (4) HiddenKiller (Qi et al., 2021b).
 The first two correspond to insertion-based methods, while the latter two involve synonym substitution and syntactic paraphrasing approaches, respectively.
+A more recent attack, BITE Yan et al.
+(2023), is discussed separately in Section˜B.3 due to its atypical behavior on benign models.
 Follow Arora et al.
@@
 The poison target labels for each dataset are listed in Table 1.
-While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section B.2.
+While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021b; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section˜B.4.
 Dataset Classes Train Test Target Class
@@
 Table 1: The statistics of our evaluated datasets.
-Defense Baselines.
-We choose six renowned defenses as our baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def.
-(He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024), (4) TIES (Yadav et al., 2023), (5) DARE (Yu et al., 2024), and (6) WAG (Arora et al., 2024).
-Among these, ONION and Z-Def.
-detect and remove triggers from datasets.
-ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def.
-detects spurious correlations between tokens and labels.
-In contrast, (2)-(6) are model purification methods.
-PURE purifies the victim model using attention head pruning and normalization techniques.
-TIES, DARE and WAG merge the victim model with other models to sanitize it.
-Evaluation Metrics.
-As introduced in  Section 3, we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure the model performance.
-CACC is evaluated on a clean test set, where both the samples and labels are ground truth.
-In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
 BadNet InsertSent LWS HiddenKiller
@@
 Z-Def. 4.6 96.1 1.8 95.6 97.3 95.3 35.7 95.4
-PURE 0 50.9 0 50.9 0 50.9 0 50.9
+PURE 0.0 50.9 0.0 50.9 0 50.9 0.0 50.9
+ABL 75.0 49.2 51.7 50.8 32.3 50.3 92.9 47.4
 TIES 99.9 95.7 100.0 95.8 93.5 95.2 88.8 95.3
@@
 PURE 2.8 86.3 3.0 85.4 6.5 85.0 9.4 84.6
+ABL 50.2 55.0 48.8 54.8 100.0 55.0 90.2 54.9
 TIES 99.9 94.6 99.6 94.4 97.7 95.8 100.0 94.4
@@
 We highlight the top-2 lowest ASR results in blue cells, and the highest CACC results in bold.
-Results for other datasets are in Table 8.
-Implementation details
+Results for other datasets are in Table˜8.
+Defense Baselines.
+We choose seven renowned defenses as main baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def.
+(He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024b), (4) ABL (Li et al., 2021b), (5) TIES (Yadav et al., 2023), (6) DARE (Yu et al., 2024), and (7) WAG (Arora et al., 2024).
+Among these, ONION and Z-Def.
+detect and remove triggers from datasets.
+ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def.
+detects spurious correlations between tokens and labels.
+In contrast, (2)-(7) are model purification methods.
+PURE purifies the victim model using attention head pruning and normalization techniques, and ABL introduces a robust anti-backdoor training framework that first isolates, then unlearns backdoor associations.
+TIES, DARE and WAG are mode-merging baselines.
+We also report three additional baselines in Section˜B.2: two model purification methods, Fine-mixing (Zhang et al., 2022) and Fine-purifying (Zhang et al., 2023), and one data filtering method, SEEP (He et al., 2024), followed by retraining.
+Evaluation Metrics.
+Following previous literature (Qi et al., 2021a), we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure utility and defense performance respectively.
+CACC is evaluated on a clean test set, where both the samples and labels are ground truth.
+In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
+Implementation details.
 Following Arora et al.
-(2024), we compare all methods on RoBERTa-large (Liu et al., 2019) and BERT-base-uncased (Devlin et al., 2019).
-All victim models are fine-tuned on poisoned datasets for three epochs using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$.
-All defense baselines are implemented based on their respective open-source repositories (see Section A.3).
-The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section 4.2.2.
+(2024), we compare all methods on RoBERTa-large (Liu et al., 2019), BERT-base-uncased (Devlin et al., 2019), as well as LLMs including Llama 2 7B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Qwen 2.5 7B (Team, 2024), with low-rank adaptation LoRA (Hu et al., 2022).
+All victim models are fine-tuned on poisoned datasets using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$.
+All defense baselines are implemented based on their open-source repositories (see Section˜A.3).
+The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section˜4.2.2.
 To train proxy models for selection, we use IMDB (Zhang et al., 2015) for the SST-2 victim model, Twitter Abusive (Founta et al., 2018) for the OLID, SNLI (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
-Our method uses hyperparameter $\alpha$ in Equation 1 to control the trade-off between backdoor defense and task utility.
+Our method uses hyperparameter $\alpha$ in Equation˜1 to control the trade-off between backdoor defense and task utility.
 By default, $\alpha$ is set to 0.4 but can be adjusted based on priorities: smaller values (e.g., 0.1) for lower ASR and larger values (e.g., 1.0) for higher utility.
-The stopping patience $T=5$ is default in Algorithm 1.
-In Section B.3, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
+The stopping patience $T=5$ is default in Algorithm˜1.
+In Section˜B.5, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
 5.2 Main results
-Table 2 presents the performance of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds.
-Results on other datasets and architectures are provided in Section B.1.
+Table˜2 presents the performance of our method compared to all baselines under benchmark backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds.
+Results on other datasets and architectures are provided in Section˜B.1.
 Across all datasets and backdoor attacks, our method consistently ranks among the top 2, regarding backdoor removal performance, with minor harm to clean accuracy.
-In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines—for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def.
+In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines–for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def.
 While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets.
-This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection.
+We attribute this to Z-Def.’s reliance on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection.
 As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures.
-For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large.
-We discuss the potential reasons for this instability in Section B.1.
-Dataset Model CACC $\uparrow$
-SST-2 Benign Victim Model 95.9
-Proxy Model (IMDB) 88.7
-Purified Model (GMS) 95.8
-AGNews Benign Victim Model 95.2
-Proxy Model (BBCNews) 78.1
-Purified Model (GMS) 95.3
-Table 3: CACC of benign victim models preserved.
+For example, it performs well on the BERT-base model in Table˜9 (following the settings reported in PURE (Zhao et al., 2024b)), but performs poorly on most tasks with RoBERTa-large.
+Unlearning-based methods (ABL Li et al.
+(2021b)) exhibit similar issues–often degrading CACC more than reducing ASR.
+Such instability has also been noted in prior works Liu et al.
+(2018); Wu et al.
+(2022), which shows that these methods are sensitive to both hyperparameter choices and the quality of the proxy data.
+Additionally, we observed that all model-merging baselines suffer in defense performance when merging with a single proxy model, aligning with the results in Arora et al.
+(2024), and tend to prioritize preserving accuracy over removing backdoors.
+We discuss the potential reasons for this and the instability of PURE in Section˜B.1.
+We also compare our approach with two earlier model purification baselines, Fine-mixing Zhang et al.
+(2022) and Fine-purifying Zhang et al.
+(2023) in Section˜B.2.
+Performance on LLMs.
+Given the growing popularity of LLMs, we evaluate the performance of our method on Llama-2-7B, Mistral-7B and Qwen-2.5-7B for the SST-2 dataset (Table˜4).
+Remarkably, in all cases, GMS effectively removes backdoors.
+[…]
+Hidden Killer 95.4 4.5
+BadNet (Diff. Trigger) 87.8 6.9
+BadNet (Same Trigger) 95.8 100.0
+Victim Model (BadNet) 95.6 100.0
+Table 3: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor is not identical in both attack strategy and trigger.
+5.3 Ablation studies
+We next examine the key property of GMS–robustness across diverse proxy datasets.
+For completeness, we defer additional results to the appendix: stability under different hyperparameter settings (Section˜B.5), transferability across attacks (Section˜B.12), preservation of clean victim model utility (Section˜B.6), and resilience to re-tuning attacks (Section˜B.10).
 BadNet InsertSent LWS HiddenKiller
-Dataset Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
+[…]
+Benign 3.6 96.7 3.5 96.7 14.9 96.7 15.2 96.7
+Victim 100.0 97.0 100.0 97.1 98.7 96.1 95.9 96.7
+Llama-2 GMS 3.5 91.7 3.6 93.4 12.2 91.5 7.6 91.1
+Benign 3.8 97.1 4.5 97.1 7.7 97.1 12.3 97.1
+Victim 100.0 95.8 100.0 96.5 98.9 95.5 95.9 96.4
+Mistral GMS 4.4 92.1 7.1 92.4 3.8 91.9 5.3 91.9
+Benign 5.0 97.0 2.7 97.0 15.1 97.0 15.7 97.0
+Victim 100.0 96.3 100.0 92.9 98.5 98.5 96.0 96.2
+Qwen-2.5 GMS 6.3 94.5 7.5 94.4 16.4 91.7 15.0 93.0
+Table 4: Performance of our method on Llama-2-7b, Mistral-7b and Qwen-2.5-7b on SST-2 dataset, averaged over three seeds.
+BadNet InsertSent LWS HiddenKiller
+Victim Model Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
 IMDB 5.4 92.8 1.6 94.0 12.6 93.8 11.0 93.0
@@
 Table 5: Proxy models trained on different datasets can all effectively purify victim models (trained on the SST-2).
-Model Dataset Backdoored Method Before Substitution After Substitution
-[…]
-Victim SST-2 BadNet 95.6 100.0 95.3 4.5
-Proxy SST-2 HiddenKiller 95.4 7.9
-Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
-Proxy SST-2 BadNet 95.6 17.1
-Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
-Proxy IMDB 87.7 8.6
-Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
-Proxy IMDB 90.1 100.0
-Table 5: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor does not exactly match in both attack strategy and trigger.
-Additionally, although merging multiple models has been shown to be an effective backdoor defense, we observed that applying all merging baselines using a single proxy model is sub-optimal, aligning with the results in Arora et al.
-(2024).
-We found these methods prioritize preserving accuracy over removing backdoors.
-This behavior likely arises from the design of merging mechanisms, which are intended to preserve the performance of each merged model on downstream tasks.
-Consequently, backdoor tasks are treated equivalently to downstream tasks, exposing a persistent vulnerability.
-5.3 Ablation studies
-Utility Preservation for Benign Victim Models
-While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign.
-Does our method compromise the utility of a clean victim model?
-As shown in Table 3, our results effectively dispel this concern.
-When applying GMS to a benign model, the CACC remains unaffected.
-The effect of different proxy models
-Our method relies on a proxy model for substitution, raising the natural question of how sensitive it is to the selection of proxy models.
-We tested two scenarios: Scenario 1: We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2.
-As shown in Table 4, our method is largely insensitive to the choice of proxy model across all backdoor attacks.
+Proxy-model robustness:
+GMS retains the robustness property of model-merging defenses: it is resilient to both the specific choice of the proxy model and the benign or malicious nature of the proxy model; i.e., as long as the backdoors in the two models are not identical, the backdoor can be effectively removed.
+We examine the sensitivity of GMS to the selection of proxy models in two scenarios.
+Scenario 1: Different proxies.
+We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2.
+As shown in Table˜5, our method is largely insensitive to the choice of proxy model across all backdoor attacks.
 Using any proxy model, GMS effectively defends against all tested attacks.
-Scenario 2: What if the proxy model itself is backdoored?
-We tested cases where the victim model and the proxy model were trained on the same dataset with different backdoor attacks or on different datasets with the same attacks.
-From Table 5, we observed that as long as the backdoor in the proxy model is not identical to that in the victim model (i.e., the same backdoor strategy with the same backdoor trigger), our GMS method consistently removes the backdoor.
-Performance under different knowledge priors
-While our method uses random sampling and outlier detection to extract the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, we also evaluate the potential performance gain if more accurate prior knowledge were available.
-For example, we may receive an accurate clean dataset or an accurate label-flipped poison dataset (i.e., dirty label) from reliable users or annotators.
-As shown in Table 6, with or without prior knowledge of accurate clean or poisoned data, our method consistently purifies the backdoor, achieving an ASR same as that of a benign model (4.1%).
-With optimal priors, GMS maintains the CACC of the benign model (95.6%), compared to 93.6% in the worst case without such prior knowledge.
-Knowledge Prior Merged SST-2 (BadNet) & IMDB
-[…]
-Random Sampling Pure Clean Outlier Detection Dirty Label
-3.2 93.6
-4.1 94.9
-4.3 92.7
-4.3 95.6
-Benign SST-2 4.1 95.6
-Table 6: Performance of purifying a victim model trained on SST-2 with a proxy model trained on IMDB, using proxy datasets with varying knowledge priors. The backdoor attack strategy is BadNet.
-BadNet InsertSent LWS HiddenKiller
-[…]
-Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
-Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
-Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
-GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
-SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
-Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
-Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
-Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
-GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
-AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
-Table 7: (Partial) The performance of transferring strategy searched based on BadNet to other attacks. Results for other datasets are in Table 12.
-Transferability of substitution strategies
-Interestingly, upon examining the module substitution strategies ($S_{best}$) searched by GMS for addressing various backdoor attacks (illustrated with examples in Figure 6 in Section B.5), we observe that, despite differences among the strategies, they exhibit transferability across different backdoor attacks.
-In Table 7, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure 6) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm 1 using $S_{badnet}$.
-Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably.
-This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.
+Scenario 2: Backdoored proxies.
+We tested cases where the victim model and the proxy model were trained on the same dataset (SST2) with different backdoor attacks (BadNet and Hidden Killer) or on different datasets (SST2 and IMDB) with the same attacks (BadNet).
+From Table˜3 and Table˜17, we observed that as long as the backdoor in proxy model is not identical to that in victim model (i.e., the same backdoor strategy with an identical backdoor trigger), GMS consistently mitigates the backdoor.
+We attribute this to the combination of module-level substitution and disruptive nature of model merging, which breaks the backdoor pathway such that a non-identical proxy cannot restore it.
+Sub-proxy datasets GMS
+Proxy Clean Proxy Poison CACC $\uparrow$ ASR $\downarrow$
+Random Sample Outlier Detection 93.6 3.2
+Random Sample Oracle 94.9 4.1
+Oracle Outlier Detection 92.7 4.3
+Oracle Oracle 95.6 4.3
+Benign SST-2 95.6 4.1
+Table 6: Performance comparison of different sub-proxy dataset configurations on GMS.
+Proxy-data robustness:
+Prior pruning- and unlearning-based methods strictly rely on access to clean data or known backdoor triggers Min et al.
+(2024); Li et al.
+(2021b); Chen et al.
+(2022); Li et al.
+(2023), typically extracted from the training subset.
+In contrast, GMS operates under more practical settings, e.g., post-processing after suboptimal data detections, where the data received is a mixture of clean and poisoned samples.
+Our previous experiments used $\mathcal{D}_{clean}$ from random sampling (containing 20% poisoned data) and $\mathcal{D}_{poison}$ from heuristic outlier detection.
+Table˜7 shows that even when 70% of the “clean” proxy dataset is poisoned, GMS is still effective in purification.
+In Appendix˜C, we mathematically justify this and discuss the analytical constraints on proxy datasets.
+These results demonstrate that our method is highly robust to proxy-data construction.
+Nevertheless, Table˜6 shows that more accurate proxy datasets–such as an oracle clean or label-flipped poison dataset–can further enhance performance, enabling precise module identification and full utility preservation (e.g., CACC of 95.6%).
+However, even without optimal proxies, our method consistently removes backdoors, achieving benign-level ASR (4.1%) with comparable CACC.
+SST-2 AGNews
+Ratio $\rho$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
+0.00 93.58 3.38 91.76 2.42
+0.30 91.51 3.15 91.59 3.61
+0.60 89.11 5.18 88.67 1.42
+0.70 89.11 5.18 88.67 1.42
+0.80 89.11 5.18 94.59 99.75
+0.90 89.11 5.18 94.83 99.82
+Table 7: Purification performance as the ratio $\rho$ of poisoned data in extracted $\mathcal{D}_{clean}$ increases.
 6 Conclusion
-We propose Greedy-Module-Substitution (GMS), an efficient and effective model purification defense against data-poisoning attacks in NLP, eliminating the need for strictly clean datasets or clean auxiliary models.
-GMS recursively identifies modules critical to backdoor pathways and substitutes them for those from a proxy model, which is not necessarily to be clean.
-Extensive experiments demonstrate the effectiveness of GMS and, more excitingly, the transferability of the optimized substitution strategies.
+We propose Guided-Module-Substitution (GMS), a model purification defense against data-poisoning attacks in NLP based on guided merging of the victim model with a single (existing) proxy model.
+Extensive experiments on both standard NLP models and LLMs validate its effectiveness, particularly against challenging attacks like LWS and HiddenKiller.
+Beyond strong empirical results, we validated desirable properties of GMS: robustness to proxy selection, tolerance to imperfect data, stability across hyperparameters, and transferability across attacks.
+These findings suggest that guided single-proxy merging offers a practical alternative to retraining-based defenses.
 Limitations
-Our method goes one step beyond previous auxiliary-model-based defenses by allowing the proxy model to be malicious, as long as its backdoor is not identical to the victim’s backdoor (i.e., the same attack method with exactly the same trigger, which is rare in practice) as shown in Table 5.
-However, further relaxation of the requirements to handle cases where the proxy model is backdoored in the same way could enable even more robust and powerful defense strategies.
-Exploring this direction presents an important and meaningful avenue for future research.
+Our method advances model-merging-based defenses by guiding the merging process with a trade-off signal between task utility and backdoor risk, reducing the requirement from a number of proxy models to just one.
+One potential limitation is that, like prior proxy-model-based defenses, we focus on the homogeneous setting and do not explore heterogeneous proxies.
+We make this choice for two reasons.
+First, in the defense scenario, merging serves solely as a tool for purification rather than an end in itself, so defender can control which proxy model to use.
+In practice, obtaining a homogeneous proxy is not difficult, making the heterogeneous case less relevant for our scenarios.
+Second, our goal is to introduce and validate the concept of guided merging for backdoor defense, rather than to develop state-of-the-art merging techniques.
+Since knowledge transfer across heterogeneous models itself remains an open challenge (Xu et al., 2025; Du et al., 2025), we leave such extensions to future work.
+Moreover, while our method tolerates the use of a malicious proxy model, as long as its backdoor is not identical to the victim’s (i.e., the same attack method with the exact same trigger, which is rare in practice), as shown in Table˜17, further relaxing this, such as enabling defenses against identically backdoored proxies or developing proxy-free approaches, could lead to even more robust and powerful defense strategies.
+Ethical statement
+Our method introduces a retraining-free purification approach for backdoored models, contributing to mitigating security risks in NLP systems from backdoor attacks.
+Additionally, our defense approach reduces the need for reannotating datasets to ensure they are purely clean, minimizing the potential harm of exposing annotators to harmfully poisoned contents.
+While we do not anticipate any direct negative consequences from this work, we hope it inspires further advancements in the development of robust, retraining-free defense methods for more realistic scenarios in future research.
+Acknowledgements
+We acknowledge the support from 2024 FSE Strategic Startup and the credits awarded from Google Cloud Platform.
 Appendix A Experiment details
@@
 We adopt the same training configuration as WAG (Arora et al., 2024) to ensure a fair comparison.
-This includes fine-tuning roberta-large and bert-base-uncased on poisoned datasets for three epochs, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$.
-The batch size and maximum sequence length are set to 32 and 128, respectively.
+This includes fine-tuning RoBERTa-large, BERT-base-uncased, as well as LLMs including Llama 2 7B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Qwen 2.5 7B (Team, 2024) on poisoned datasets, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$.
+For encoder models, the batch size, maximum sequence length, and epoch are set to 32, 128, and 3, respectively.
+For decoder LLMs, due to computational limits, we adapt low-rank adaptation (LoRA) (Hu et al., 2022) for all the linear modules within the Transformer layers and apply our method to substitute those LoRA modules after training.
+We train these LLMs for 2 epochs with a batch size of 8 and a maximum sequence length of 128.
 To validate our defense, we first adapt the aforementioned backdoor methods and datasets to obtain victim models.
+Specifically, to train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively.
+For MNLI, we use a random subset of 100,000 samples to reduce overhead.
 Next, we train proxy models on homologous datasets to prepare for future merging.
-We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section 4.2.2) to guide our parameter substitution strategy search.
+We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section˜4.2.2) to guide our parameter substitution strategy search.
 For a model trained on unverified data, access to fully clean data is often not guaranteed.
@@
 Specifically, we use Twitter Abusive (Founta et al., 2018) for OLID, SNLI [^3] Creative Commons Attribution-ShareAlike 4.0 International License (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
-To evaluate generalizability, we use three sentiment classification datasets—IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)—to purify victim SST-2 models through model merging.
+To evaluate generalizability, we use three sentiment classification datasets–IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)–to purify victim SST-2 models through model merging.
 All datasets are downloaded from Hugging Face, which adheres to the Apache License 2.0.
-Our method involves only one hyperparameter, $\alpha$, as described in  Equation 1, which controls the preference between model utility and attack resistance during the model merging strategy search.
+Our method involves only one hyperparameter, $\alpha$, as described in  Equation˜1, which controls the preference between model utility and attack resistance during the model merging strategy search.
 We set $\alpha$ to 0.4 by default, which slightly favors seeking a more attack-resistant model.
@@
 A.2 Computational Resources
-We conduct experiments using three seeds on Nvidia V100 and A1000 GPUs, and report the average scores. Running our method takes only 4 minutes on a single GPU.
+We conduct experiments using three seeds on a single A100 GPU, and report the average scores. Running our method takes only 4 minutes on a single GPU for a 24-layer RoBERTa-large architecture.
 BadNet InsertSent LWS HiddenKiller
@@
 We retained the original search space for drop rate and scaling term but expanded the ratio to retain from [0.1, 0.2, 0.3] to [0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8] for the roberta-large model to ensure more complete search.
-For the baseline PURE (Zhao et al., 2024), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts.
+For the baseline PURE (Zhao et al., 2024b), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts.
 The pruned model is then fine-tuned further on the proxy clean dataset to normalize the remaining attention heads for purification.
@@
 Performance across different dataset.
-Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table 2.
-The settings are consistent with those in Section 5.2, with each score averaged over three runs using different seeds.
-For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure 4.
+Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table˜2.
+The settings are consistent with those in Section˜5.2, with each score averaged over three runs using different seeds.
+For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure˜3.
 The top two ASR performances are highlighted.
@@
 As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures.
-For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large.
+For example, it performs well on the BERT-base model in Table˜9 (following the settings reported in PURE (Zhao et al., 2024b)), but performs poorly on most tasks with RoBERTa-large.
 This instability seems to arise from PURE’s reliance on accuracy from the clean proxy dataset as the pruning stopping criterion.
 For well-trained models with high accuracy, a substantial number of heads are pruned, leading to a broken purified model (e.g., 0% ASR but random-guess-level CACC).
+Additionally, although merging multiple models has been shown to be an effective backdoor defense, we observed that applying all merging baselines using a single proxy model is sub-optimal, aligning with the results in Arora et al.
+(2024).
+We found these methods prioritize preserving accuracy over removing backdoors.
+This behavior likely arises from the design of merging mechanisms, which are intended to preserve the performance of each merged model on downstream tasks.
+Consequently, backdoor tasks are treated equivalently to downstream tasks, exposing a persistent vulnerability.
 Performance across different model architectures.
 We also evaluated our method on different architectures.
-Table 9 shows the results on the BERT-base model.
+Table˜9 shows the results on the BERT-base model.
 The poison rate remains at 20%, and each model is trained for three epochs.
@@
 Notably, it achieves over a 39% advantage in defending against LWS and a 9% improvement over the second-best baseline for HiddenKiller.
-B.2 Performance under different poison rate: 20%, 10% and 5%
-While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table 10.
+B.2 Comparisons to other baseline defenses
+Model purification baselines Fine-mixing Zhang et al.
+(2022) and Fine-purification Zhang et al.
+(2023) are two additional model purification baselines.
+We compare our method against their reported performance in their papers [^4] The code is not publicly available yet., strictly following their experimental settings.
+Specifically, we evaluate under the same conditions using two types of attacks (BadNet and InsertSent) and two model architectures (BERT-base and RoBERTa-base) on the AGNews dataset.
+As shown in Table˜10, our method consistently achieves better defense performance across all settings.
+Notably, against the more advanced InsertSent attack, our approach reduces the attack success rate (ASR) by at least 17% in absolute values compared to Fine-mixing and Fine-purification.
+Architecture Methods BadNet InsertSent
+CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
+BERT-base Fine-mixing 90.17 12.32 90.40 32.37
+Fine-purifying 90.86 3.3 91.13 23.69
+Ours 91.8 1.8 86.49 7.00
+RoBERTa-base Fine-mixing 86.39 18.12 86.11 35.97
+Fine-purifying 86.64 17.56 86.85 19.20
+Ours 91.2 1.63 89.62 1.42
+Table 10: Performance comparison of our method against Fine-mixing Zhang et al.
+(2022) and Fine-purifying Zhang et al.
+(2023) on AGNews dataset.
+The best results are in bold.
+Results in rows marked with are taken directly from the respective papers.
+Poison sample detection with retraining: SEEP He et al. (2024)
+While this paper focuses on the important but often overlooked post-deployment setting, it is natural to ask: if we had access to a highly effective data detection method capable of identifying most poisoned samples, how would our method compare?
+To explore this, we evaluate against SEEP He et al.
+(2024), a state-of-the-art poison detection method, on defending against BadNet attacks on the SST-2 dataset.
+As shown in Table˜11, although SEEP performs well under this relatively simple attack, our method–despite relying only on a naive heuristic for proxy data selection–achieves even better performance.
+As further discussed in Table˜6, incorporating more advanced prior knowledge (e.g., by using stronger detection methods to construct sub-proxy datasets) can further enhance the effectiveness of our approach.
+[…]
+RoBERTa-large SEEP with retraining 95.5 24.3
+Ours 91.7 9.7
+BERT-base SEEP with retraining 92.4 29.4
+Ours 88.0 27.4
+Table 11: Comparison of our method with retraining-based defenses on BERT-base and RoBERTa-large for SST-2 dataset.
+B.3 Comparisons of all defenses under a more recent attack: BITE Yan et al. (2023)
+BITE Yan et al.
+(2023) is a recent insertion-based textual backdoor attack that leverages label-biased tokens as stealthy triggers.
+Table˜12 shows the performance of our method and four baselines in defending against BITE on the SST-2 dataset.
+Our method consistently ranks among the top two in defense effectiveness, achieving comparable or even lower ASR than the benign model, while maintaining clean accuracy.
+We exclude BITE from the main results in Section˜5.2 due to a peculiar issue: BITE tends to yield an unusually high ASR even on benign models.
+This undermines the reliability of ASR as an evaluation metric in this setting and makes it difficult to confidently interpret defense performance.
+We attribute this to BITE’s trigger selection process, which biases trigger tokens toward naturally label-correlated words, effectively exploiting model priors without requiring explicit poisoning.
+Model Defense Method BERT-base RoBERTa-Large
+[…]
+Victim (BITE) Benign Model 92.9 41.4 95.9 38.3
+No Defense 92.6 80.7 95.6 81.0
+Defense Z-Defense 92.3 51.9 95.2 44.9
+ONION 89.7 70.5 92.9 68.2
+ABL 92.0 82.1 51.6 1.0
+WAG 92.9 47.5 94.4 45.2
+Ours 88.9 49.9 90.3 27.9
+Table 12: Defense performance against the BITE Yan et al. (2023) attack on BERT-base and RoBERTa-large models for SST-2 dataset.
+B.4 Performance under different poison rate: 20%, 10% and 5%
+While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021b; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table˜13.
 Our method is effective in model purification across all poison rates.
@@
 Figure 3: Results of using different weights (alpha) on AGNEWS dataset.
-B.3 How sensitive is our method to the selection of hyperparameters?
-Our method involves two hyperparameters: $\alpha$ in Equation 1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results.
+B.5 Stability: How sensitive is our method to the selection of hyperparameters?
+We next show that GMS exhibits stable performance across a wide range of hyperparameters, in contrast to pruning-based and unlearning-based methods, which have been shown to be more sensitive to hyperparameter choices Liu et al.
+(2018); Wu et al.
+(2022) (as also observed in Table˜2).
+Our method involves two hyperparameters: $\alpha$ in Equation˜1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results.
 Below, we justify these choices.
-As shown in Figure 4 and Figure 4, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength.
+As shown in Figure˜3 and Figure˜3, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength.
 An exception is the HiddenKiller attack on the OLID dataset, where a more aggressive $\alpha=0.1$ is recommended.
@@
 This change significantly impacts CACC, necessitating a higher weight for backdoor removal to ensure effective purification.
-For stopping patience, as shown in Figure 5, the score computed in Equation 1 generally follows a (weakly) monotonically decreasing trend.
+For stopping patience, as shown in Figure˜4, the score computed in Equation˜1 generally follows a (weakly) monotonically decreasing trend.
 While a larger $T$ improves performance, we choose $T=5$ as a reasonable balance between performance and efficiency.
 Figure 4: The search iteration history for four kinds of backdoors on SST-2 dataset.
+Dataset Model CACC $\uparrow$
+SST-2 Benign Victim Model 95.9
+Proxy Model (IMDB) 88.7
+Purified Model (GMS) 95.8
+AGNews Benign Victim Model 95.2
+Proxy Model (BBCNews) 78.1
+Purified Model (GMS) 95.3
+Table 14: CACC of benign victim models preserved.
+[…]
+Pure clean proxy dataset 88.3 4.7
+Mixed poisoned proxy dataset 91.6 4.5
+Table 15: Performance of our method with a clean vs. mixed proxy dataset, for the roberta-large model trained on the SST-2 dataset.
 Methods
@@
 Figure 5: The optimal substitution strategy for defending against each backdoor attack for the roberta-large model trained on the SST-2 dataset. The green squares indicate the substituted modules.
-B.4 Robustness to Retuning Attacks
+B.6 If the victim model is benign, will our method affect its utility?
+While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign.
+Does our method compromise the utility of a clean victim model?
+As shown in Table˜14, our results effectively dispel this concern.
+When applying GMS to a benign model, the CACC remains unaffected.
+B.7 Can our method be effective with a fully clean proxy dataset?
+While we primarily consider the practical scenario of receiving a mixed dataset, is our method still effective if the received dataset is entirely clean?
+As shown in Table˜15, when provided only with a clean proxy dataset—without any information about the poisoned dataset—GMS can still effectively remove backdoors.
+This is because our substitution strategy starts by fully removing backdoors and then gradually recovering task utility.
+B.8 GMS remains effective as long as the proxy model is not backdoored with exactly the same attack and trigger.
+Model Dataset Backdoored Method Before Substitution After Substitution
+[…]
+Victim SST-2 BadNet 95.6 100.0 95.3 4.5
+Proxy SST-2 HiddenKiller 95.4 7.9
+Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
+Proxy SST-2 BadNet 95.6 17.1
+Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
+Proxy IMDB 87.7 8.6
+Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
+Proxy IMDB 90.1 100.0
+Table 17: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor does not exactly match in both attack strategy and trigger.
+An important question to consider is: What happens if the proxy model itself contains a backdoor?
+To investigate this, we considered scenarios where the victim and proxy models were either trained on the same dataset but with different backdoor attacks or on different datasets with the same attack.
+From Table˜17, our findings indicate that GMS remains effective in backdoor removal as long as the proxy model’s backdoor is not an exact match to that of the victim model, i.e., both the same attack method and the same trigger.
+B.9 Purification effectiveness as the poison ratio increases.
+Due to space constraints, we included an incomplete version of Table˜7 in the main paper.
+Here, the full ratio table Table˜18 provides a clearer trend: (1) As the proportion of poisoned data in the proxy “clean” dataset increases, the CACC of the purified model slightly decreases.
+This suggests that under a default weight of $\alpha=0.4$ for $\mathcal{D}_{clean}$, a higher level of poisoning may introduce confusion, leading GMS to mistakenly remove some clean task-critical modules due to their reduced relative weight (see Appendix˜C for the analytical explanation).
+(2) However, even with poisoned ratios as high as 90% in SST-2 and 70% in AGNews, GMS remains effective in eliminating backdoor-related components and purifying the model.
+This effectiveness is mathematically justified in Appendix˜C.
+SST-2 AGNews
+Ratio $\rho$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
+0.00 93.58 3.38 91.76 2.42
+0.10 93.58 3.38 91.76 2.42
+0.20 91.06 4.28 91.59 3.61
+0.30 91.51 3.15 91.59 3.61
+0.40 89.11 5.18 91.59 3.61
+0.50 89.11 5.18 91.59 3.61
+0.60 89.11 5.18 88.67 1.42
+0.70 89.11 5.18 88.67 1.42
+0.80 89.11 5.18 94.59 99.75
+0.90 89.11 5.18 94.83 99.82
+Table 18: Purification performance as the ratio $\rho$ of poisoned data in the proxy “clean” dataset $\mathcal{D}_{clean}$ increases.
+B.10 Robustness to Retuning Attacks
 Recent works Min et al.
@@
 (2024) have highlighted a critical pitfall in backdoor defenses: purified models with low ASR often fail to completely eliminate inserted backdoor features.
-A straightforward Retuning Attack (RA)—which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)—can easily recover the backdoor in the purified model, as demonstrated in Table 11.
+A straightforward Retuning Attack (RA)–which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)–can easily recover the backdoor in the purified model, as demonstrated in Table˜16.
 Several related studies Qi et al.
@@
 This challenge motivates us to propose a purification method aimed at breaking this backdoor-connected path.
-Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024) or merging them with other parameters Arora et al.
+Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024b) or merging them with other parameters Arora et al.
 (2024); Yadav et al.
@@
 Moreover, because parameters often interact within functional modules, modifying at the module level rather than the individual parameter level tends to remove backdoor features more effectively.
-As shown in Table 11, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al.
+As shown in Table˜16, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al.
 (2024), the parameter-merging baseline, WAG Arora et al.
@@
 While the recent method PURE also demonstrates competitive performance, it is worth noting that significant effort was required to find a configuration where its pruning step did not break the purified model.
-Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%—nearly twice the robustness.
-B.5 Examples of Searched Substitution Strategy
+Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%–nearly twice the robustness.
+B.11 Examples of Searched Strategy
 We applied our GMS defense to purify victim models compromised by various backdoor attacks across four datasets.
 Examples of the searched strategies from the SST-2 dataset are shown in Figure 5.
-While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section 5.3 and  Section B.6.
+While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section˜B.12.
 BadNet InsertSent LWS HiddenKiller
@@
 Table 19: The performance of transferring strategy searched based on BadNet to other attacks.
-B.6 Transferability of substitution strategies
-We present the results for SST-2 and AGNews in Table 7, with additional results for other datasets provided in Table 12.
-All experiments follow a consistent setup: the substitution strategy is optimized based on BadNet attacks and then applied to coordinate module substitutions for other attacks.
-For instance, in the MNLI task, the substitution strategy identified for defending against BadNet—using a proxy model trained on SNLI—is directly applied to replace modules in InsertSent, LWS, and HiddenKiller with those from the proxy model.
-Across all tasks, the defense performance is surprisingly strong, revealing an intriguing observation: a substantial overlap exists in the components impacted by different attacks.
+B.12 Transferability of substitution strategies
+Interestingly, we find that GMS exhibits strong transferability: a substitution strategy optimized for one attack often generalizes to others, enabling defense even without access to data.
+Visually, the module substitution strategies ($S_{best}$) identified by GMS across different backdoor attacks (examples in Figure˜5 in Section˜B.11), though distinct, share many similar patterns across attacks.
+In Table˜19, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure˜5) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm˜1 using $S_{badnet}$.
+Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably.
+This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.
+Appendix C Constraints on the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
+Table˜7 and Table˜6 illustrates that our method is quite robust to the construction of the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, e.g., even random sampling (for $\mathcal{D}_{clean}$) and inaccurate heuristics (for $\mathcal{D}_{poison}$) can yield effective defense.
+Natural questions arise: Are there any constraints on the proxy datasets?
+Could the random sampling strategy for constructing $\mathcal{D}_{clean}$ fail if $\mathcal{D}_{poison}$ is highly inaccurate?
+We next establish the constraints and demonstrate that, in most cases, even when only half of the extracted samples in $\mathcal{D}_{poison}$ are poisoned, our approach can still use random sampling for $\mathcal{D}_{clean}$ to effectively purify the model.
+Let the constant values $s_{asr}(M_{v})$ and $s_{acc}(M_{v})$ in Equation˜2 and Equation˜3 be denoted as $c_{1}$ and $c_{2}$, respectively, for simplicity.
+We use ASR and ACC to represent the ground-truth attack success rate and clean accuracy of a model (that can be measured on oracle poisoned and clean test sets), respectively.
+Assuming we have two imperfect proxy datasets: a proxy “clean” dataset that contains a proportion $\rho$ of poisoned data and a proxy “poisoned” dataset contains a proportion $1-\lambda$ of the non-poisoned data.
+Denote the poisoned and clean data distributions as $\mathcal{P}$ and $\mathcal{C}$, respectively.
+Then, by the rule of total probability and definition in Section˜3, we obtain:
+$s_{acc}(M)$ $=\mathbb{E}_{x\sim\mathcal{D}_{clean}}[\mathbf{1}(M(x)=y)]$
+$=p(x\in\mathcal{P}\mid x\sim\mathcal{D}_{clean})$
+$\quad\cdot p(M(x)=y\mid x\in\mathcal{P},x\sim\mathcal{D}_{clean})$
+$\quad+p(x\in\mathcal{C}\mid x\sim\mathcal{D}_{clean})$
+$\quad\cdot p(M(x)=y\mid x\in\mathcal{C},x\sim\mathcal{D}_{clean})$
+$=\rho\text{ASR}+(1-\rho)\text{ACC}.$
+Similarly, for $x\sim\mathcal{D}_{poison}$, we have:
+$s_{asr}(M)$ $=\lambda\cdot\text{ASR}+(1-\lambda)\cdot\text{ACC}.$
+Substituting these into Equation˜1 gives us $(1-\alpha)c_{1}+\alpha(1-c_{2})+\left(\alpha\rho+\alpha\lambda-\lambda\right)\text{ASR}+\left(2\alpha-1-\alpha\rho-\alpha\lambda+\lambda\right)\text{ACC}$.
+From this, the constraints ensuring effective purification are:
+$\left\{\begin{array}[]{ll}\alpha\rho<(1-\alpha)\lambda,\\ \alpha(1-\rho)>(1-\alpha)(1-\lambda).\end{array}\right$. (6)
+That is, as long as the clean portion in $\mathcal{D}_{clean}$ receives more attention than that in $\mathcal{D}_{poison}$, our method will effectively remove backdoor modules while preserving utility; the same applies to the poisoned portion in reverse.
+To further illustrate, for an accurate proxy dataset, the constraints are simplified to $\alpha(1+\rho)-1<0$ and $\alpha(1-\rho)>0$.
+Thus, if $\rho<1$, setting $\alpha$ such that $0<\alpha<\frac{1}{\rho+1}$ can ensure effective purification.
+In our experiments, $\alpha$ is set to 0.4 by default.
+Then, the requirement becomes $\rho<\frac{3}{2}\lambda-\frac{1}{2}$.
+This implies that even with a random-guess backdoor data detection method, we can tolerate up to $25\%$ poisoned data in $\mathcal{D}_{clean}$.
+Given that real-world poisoning rates are usually low (e.g., below $20\%$ or even $1\%$), random sampling for the “clean” proxy dataset remains highly feasible in practice.
+Appendix D AI Assistants
+We use ChatGPT/Gemini for writing and formatting supports, including grammar checks, improving the clarity of figure and table captions, and other surface-level edits.